### v1

# Chaos Engineering: A Multi-Vocal Literature Review

Abstract.
Organizations, particularly medium and large enterprises, typically today rely heavily on complex, distributed systems to deliver critical services and products. However, the growing complexity of these systems poses challenges in ensuring service availability, performance, and reliability. Traditional resilience testing methods often fail to capture modern systems’ intricate interactions and failure modes. Chaos Engineering addresses these challenges by proactively testing how systems in production behave under turbulent conditions, allowing developers to uncover and resolve potential issues before they escalate into outages. Though chaos engineering has received growing attention from researchers and practitioners alike, we observed a lack of a comprehensive literature review. Hence, we performed a Multivocal Literature Review (MLR) on chaos engineering to fill this research gap by systematically analyzing 88 academic and grey literature sources published from January 2019 to April 2024. We first used the selected sources to derive a unified definition of chaos engineering and to identify key capabilities, components, and adoption drivers. We also developed a taxonomy for chaos engineering and compared the relevant tools using it. Finally, we analyzed the state of the current chaos engineering research and identified several open research issues.
Chaos Engineering, Fault Injection, Distributed System, Resilience, Multivocal Literature Review, Microservice
copyright: noneisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software testing and debugging
1. Introduction
Cloud-based applications are being adopted across various domains, including healthcare, finance, construction, education, and defense (Alabbadi, 2011; Dang et al., 2019; Bello et al., 2021). While cloud computing promises to offer benefits such as cost savings, flexibility, and faster time to market (Buyya et al., 2009), the growing complexity of cloud infrastructure poses a significant challenge to maintaining dependable systems, leading to an increased risk of technical failures and costly downtime (Song and Sohn, 2022; Li et al., 2011; Wang et al., 2015; Al-Said Ahmad et al., 2024). The consequences of these failures can be severe, as seen in significant financial losses incurred for companies, including the 2007-2013 outages of 38 cloud service providers, resulting in \$669 million in losses, and Facebook’s six-hour outage in 2021, which resulted in an estimated \$160 million loss for the global economy  (Gill, 2022). The 2017 ITIC (Information Technology Intelligence Consulting) survey also reported that 98% of organizations experienced downtime with financial losses exceeding \$100,000 per hour (Inc., 2023).
An effective approach to overcoming the challenges of ensuring reliable and resilient cloud-based systems is chaos engineering, which proactively tests system resilience by introducing controlled failures into production environments (Bhaskar, 2022; Fawcett, 2020; Palani and Gupta, 2023). This approach was pioneered by Netflix in 2008 to fortify its video streaming infrastructure. Since then, chaos engineering has grown and extended its influence, evolving from its original focus on cloud services and microservice architecture to broader applications across industries (Bairyev, 2023). Today, this approach is widely employed in critical sectors like fintech and healthcare, with major technology companies such as LinkedIn, Facebook, Google, Microsoft, Amazon, Slack, and others adopting it to enhance their system resilience, improve customer experience, and drive business growth.  (Inc., 2023; Patel, 2022; Roa, 2022; Fawcett, 2020; Al-Said Ahmad et al., 2024; Bremmers, 2021; Bairyev, 2023).
Figure 1 illustrates chaos engineering’s adoption through Moore’s ’Crossing the Chasm’ model (Moore and McKenna, 1999). Adoption begins with Innovators (early risk-takers) and Early Adopters (those seeking competitive advantage), followed by the Early Majority (who adopt once proven), the Late Majority (who wait until widely supported), and finally, Laggards (the last to adopt, often only when essential). The critical hurdle, Moore’s Chasm, is where many approaches fail due to complexity and unclear benefits. Crossing this chasm leads to mainstream adoption and integration across the SDLC (Software Development Lifecycle) and industries (Mukkara, 2022).
Figure 1. Chaos Engineering in the Context of Geoffrey Moore’s Technology Adoption Lifecycle Model from Crossing the Chasm  (Mukkara, 2022).
[…]
Chaos Engineering on the Technology Adoption Lifecycle
We observed many researchers and practitioners have explored chaos engineering in various contexts, discussing its implementations and challenges through research papers and grey literature such as online blogs, technical papers, and forums. However, despite its increasing adoption, much of the existing knowledge remains dispersed across the literature. To the authors’ best understanding, no secondary research has been published to synthesize this scattered information, such as comprehensive reviews or survey papers. Hence, to address this research gap, in this paper, we present a Multivocal Literature Review (MLR) (Garousi et al., 2019) on chaos engineering, synthesizing a broad spectrum of perspectives from academic and grey literature sources. Unlike traditional Systematic Literature Reviews (SLRs), which often overlook grey literature, an MLR includes insights from peer-reviewed academic studies alongside industry reports, technical blogs, and white papers (Garousi et al., 2019, 2016). In particular, as chaos engineering emerged from the industry and most tools are from the industry, the grey literature is a highly relevant and valuable source of information.
In this MLR, we selected and analyzed 88 sources, spanning academic and grey literature sources. Our findings provide a detailed overview of chaos engineering, including its core principles, quality requirements, and foundational components defined by researchers and industry experts. We also examine the motivations behind adopting chaos engineering, the specific resilience challenges it addresses, and the practices and evaluation methods that support its effective implementation. Additionally, we introduce a taxonomy that can help practitioners and decision-makers make informed decisions when comparing and selecting chaos engineering tools. By analyzing the benefits and shortcomings of current approaches in the chaos engineering domain, this MLR contributes to a thorough understanding of chaos engineering and provides guidelines for future studies to advance this critical field.
The remainder of the paper is organized as follows. Section  2 presents the research methodology used in this study, while Section 3 provides a comprehensive definition of chaos engineering, along with its functionalities, quality requirements, and core components. Section 4 focuses on the motivations for chaos engineering and the specific challenges it addresses. Section 5 introduces a taxonomy of chaos engineering, covering tools, practices, and evaluation methods within the field. Section 6 examines the current state of chaos engineering in research and industry. Section  7 discusses the implications of our findings, highlighting key insights, identifying open issues, and addressing potential threats to validity. Finally, Section 8 summarizes the related literature surveys, and Section  9 concludes the paper.
2. Research Methodology
We employed the guidelines proposed by Garousi et al. (Garousi et al., 2019) to systematically collect and analyze multivocal literature on chaos engineering. We also drew methodological inspiration from other multivocal literature reviews, such as Buck et al. (Buck et al., 2021) and Islam et al. (Islam et al., 2019). Figure  2 illustrates our MLR methodology, comprising several steps divided into three phases: planning and designing the review protocol, executing the review, and documenting the findings. The remainder of this section provides a detailed discussion of each step.
2.1. Research Identification
We employed a search strategy to identify appropriate literature grounded in a series of research questions in Table  1. The first three questions cover the various aspects of chaos engineering, including its definitions, key features, benefits, challenges, use cases, tools and techniques, and best and bad practices. The final question provides an overview of the present landscape of chaos engineering literature.
Figure 2. Overview of the MLR Process(Islam et al., 2019).
[…]
Multivocal Literature Review Process.
Table 1. Research Questions and Motivation
Research Questions Motivation
RQ1. How is chaos engineering defined in the literature? RQ1.1. Which core principles and primary functionalities are most commonly associated with chaos engineering? RQ1.2. What are the core components of a chaos engineering platform, and how do they support its core functions? RQ1.3. What quality requirements are essential for an effective chaos engineering platform, and how are they quantified? This question clarifies the definition of chaos engineering, exploring its core principles, key components, and quality requirements. These elements build a foundation for understanding, implementing, and evaluating chaos engineering solutions.
RQ2. What challenges does chaos engineering intend to solve? This question helps identify key challenges, guiding organizations to areas where chaos engineering can best enhance resilience and reliability.
RQ3. What kinds of solutions have been proposed for chaos engineering? RQ3.1. What tools and techniques are utilized, proposed, designed, and implemented by researchers and practitioners? RQ3.2. What best and bad practices are documented in adopting chaos engineering? RQ3.3. How are the chaos engineering approaches evaluated? This question examines the tools and techniques in chaos engineering, highlighting key practices, common pitfalls, and methods for evaluating and enhancing solutions across various contexts.
RQ4. What is the state of chaos engineering in research and industry? RQ4.1. What is the extent and diversity of research conducted on chaos engineering, and what trends can be observed over time? RQ4.2. Who are the thought leaders and key contributors in the field of chaos engineering, and what intersections of thought do they represent? Exploring this question allows us to assess the growth and direction of chaos engineering, highlighting current trends, influential contributors, and areas that need further exploration, ensuring continued advancement in the field.
2.2. Search Strategy
This section outlines the search strategy, detailing the data sources, queries, and the complete search process.
2.2.1. Sources of data
We conducted automated searches in four major digital libraries—Scopus, IEEE Xplore, ACM Digital Library, and Web of Science—using both peer-reviewed scholarly articles and grey literature as sources. These libraries serve as primary sources for software engineering studies (Chen et al., 2010). We performed multi-keyword searches targeting titles, abstracts, and keywords for papers published between January 2019 and April 2024. Additionally, Google Scholar was used to pinpoint any relevant literature that was not included in the previous searches. We used Google Search for grey literature, similar to other multi-vocal literature reviews (MLRs)  (Garousi and Mäntylä, 2016; Garousi et al., 2017; Islam et al., 2019).
2.2.2. Search Terms
Following the SLR guidelines (Kitchenham et al., 2007), we constructed search terms based on the research questions. The first two authors tested the search strings to assess their applicability for addressing the defined research questions. Multiple pilot searches were conducted to refine the search terms, ensuring the inclusion of all relevant papers—boolean operators (AND, OR) combined keywords, synonyms, and related concepts. Table 4 shows the final search strings used across the four digital libraries. We adapted and applied the Google Search search string “(Chaos AND Engineering)” for grey literature to capture relevant non-peer-reviewed sources.
2.3. Eligibility Criteria
Following some existing MLR studies (Islam et al., 2019; Butijn et al., 2020; Garousi et al., 2019), we applied a specific set of inclusion and exclusion criteria, as detailed in Table  2
Table 2. Inclusion and Exclusion Criteria.
Inclusion Criteria Exclusion Criteria
IC1: Articles must be in English and be readily accessible in full text. EC1: Short articles or brief papers (less than six pages).
IC2: Articles must include robust validation methods, including working prototypes or tools with proper references, especially for grey sources. EC2: Articles that lack the relevant focus on chaos engineering.
IC3: Articles that report practices and challenges in implementing chaos engineering, including case studies, best practices, or guidelines. EC3: Articles that predominantly discuss physical infrastructure or purely hardware aspects without direct connection to chaos engineering.
IC4: Articles that provide insights into the trends and evolution of chaos engineering practice. EC4: Studies enhancing specific algorithms or features of a single tool without a broader chaos engineering context unless they provide insights into the field’s evolution or state of tools.
IC5: Articles that identify key contributors and diverse perspectives within chaos engineering. EC5: Duplicate articles from all sources.
IC6: Articles must cover all identified phases of the chaos engineering pipeline, ensuring a holistic view of the field. EC6: Articles lacking empirical evidence or practical applications to support theoretical propositions.
2.4. Study Selection
The libraries’ advanced query modes were used to maximize search coverage. The search was conducted in April 2024, and the results were saved in BibTeX and CSV formats. Duplicates were removed before proceeding with further analysis. Figure  3 illustrates the selection process for academic and grey literature, showing the databases searched and the number of selected papers at each stage. Different approaches were followed to select grey and academic literature.
2.4.1. Academic Literature Selection
Searches in Scopus, IEEE Xplore, ACM DL, and Web of Science yielded 1,290 papers. After duplicate removal, 222 unique papers remained. Applying the inclusion and exclusion criteria, the selection was narrowed to 42 relevant papers. An additional search was conducted using Google Scholar, which added two more papers to the final count, bringing the total to 44 papers.
Figure 3. MLR Study Selection Process.
[…]
MULTIVOCAL_REVIEW_Final
Table 3. Criteria for Qualitative Assessment of Grey Literature.
Criteria Questions
Authority (Q1). Is the organization publishing the work credible?
(Q2). Are there other publications by the author in this domain of study?
(Q3). How closely is the author connected to a reputable institution?
Methodology (Q4). Is the objective of the source evident and well-stated?
(Q5). Can the author be considered an expert in this field?
Date (Q6). Is the publication date of the source within the past five years?
Table 4. Search Strings and Results.
Search Strings ACM DL IEEE Xplore Scopus Web of Science Total
"chaos engineering" OR "chaos test" OR "chaos experiment" OR "chaos toolkit" OR "chaos platform" OR "chaos mesh" OR "chaos integration" OR "chaos architecture" OR "chaos infrastructure" OR "chaos injection" OR "chaos tool" 95 63 121 47 326
"chaos engineering" AND ("trends" OR "developments" OR "progression") 29 0 17 0 46
"chaos engineering" AND ("leader" OR "pioneer" OR "expert" OR "contributor") 35 3 4 0 42
"chaos engineering" AND ("phases" OR "stages" OR "steps" OR"pipeline") 62 5 11 0 78
"chaos engineering" AND ("challenges" OR "issues" OR "problems") 79 16 34 4 133
"chaos engineering" AND ("applications" OR "benefits") 77 20 37 8 142
"chaos engineering" AND ("solutions" OR "approaches" OR "methods") 72 20 55 4 151
"chaos engineering" AND ("tools" OR "techniques") 79 24 36 8 147
"chaos engineering" AND ("best practices" OR "bad practices" OR "implementation" OR "Culture") 68 8 16 5 97
"chaos engineering" AND ("evaluation" OR "metrics") 75 24 26 3 128
Total 671 183 357 79 1290
2.4.2. Grey Literature Selection
We followed the approach of other multi-vocal and grey literature reviews  (Garousi and Mäntylä, 2016; Garousi et al., 2017; Islam et al., 2019; Butijn et al., 2020; Garousi et al., 2019) and used Google Search to gather grey literature. As described by these reviews, we examined the first ten pages of results, as this was deemed sufficient to identify the most pertinent literature since Google’s algorithm emphasizes the most relevant results on the first pages  (Soldani et al., 2018; Verdecchia et al., 2019; Islam et al., 2019). We continued scanning further only if necessary, stopping when no relevant new articles appeared  (Islam et al., 2019). To avoid bias, we searched in incognito mode, following the recommendations of Rahman et al. (Rahman et al., 2023). We incorporated additional quality assessment criteria beyond the inclusion and exclusion criteria, as outlined in Table  3. We picked six key criteria: the credibility of the publishing organization (Q1), the author’s prior publications in the field (Q2), the author’s institutional affiliation (Q3), clarity of purpose (Q4), the author’s expertise (Q5), and recency of the source (Q6). The responses to these criteria were classified as ‘Yes’ (1), ‘Partly’ (0.5), and ‘No’ (0), with scores assigned accordingly for each category. These criteria were adapted from prior studies on grey and multi-vocal literature reviews  (Soldani et al., 2018; Butijn et al., 2020; Kumara et al., 2021), ensuring a thorough evaluation of the sources. As a result, we selected 44 documents, contributing to 88 unique articles (Figure  3) for extracting and synthesizing data. Table  5 illustrates the selected sources’ distribution across various venues.
Table 5. Selected Study for Data Synthesis and Qualitative Evaluation.
Academic Literature Grey Literature
Websites & Blogs Whitepaper
(Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Zhang et al., 2021a; Cotroneo et al., 2022; Camacho et al., 2022; Torkura et al., 2021; Poltronieri et al., 2022; Zhang et al., 2023; Nikolaidis et al., 2021; Frank et al., 2021b, a; Pierce et al., 2021; Soldani and Brogi, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Chen et al., 2024; Nikolaidis et al., 2023; Ma et al., 2023; Meiklejohn et al., 2021; Frank et al., 2023; Sondhi et al., 2021; Kassab et al., 2022; Jernberg et al., 2020; Konstantinou et al., 2021; Pulcinelli et al., 2023; Ji et al., 2023; Ikeuchi et al., 2020; Chen et al., 2022; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Basiri et al., 2019; Park et al., 2019; Klinaku et al., 2022; Rivera et al., 2023; Camilli et al., 2022; Yu et al., 2021; Siwach et al., 2022) = 44 (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Patel, 2022; Jakkaraju, 2020; Hawkins, 2020; Nombela, 2023; Le, 2022; National Australia Bank, 2020; Shah, 2021; Krishnamurthy, 2021; Moon, 2022; Katirtzis, 2022; Wachtel, 2023; Durai, 2022; Anwar and Arunachalam, 2019; Tomka, 2024; Dua, 2024; Bocetta, 2019; Mooney, 2023; Wu, 2021; Lardo, 2019; Miles, 2019a; Bhaskar, 2022; Azure Readiness, 2023; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Bairyev, 2023; Sbiai et al., 2023; Abdul, 2024; Roa, 2022; Kadikar, 2023; Green, 2023; Aharon, 2024; Krivas and Portela, 2020; Kostic, 2024; Vinisky, 2024; Colyer, 2019) = 41 (Palani and Gupta, 2023; Tivey et al., 2019; Starr et al., 2022) = 3
2.5. Data Extraction, Synthesis, and Analysis
In alignment with the MLR process, we read and assessed the selected papers, extracting and summarizing data to address our research questions.
2.5.1. Data Extraction
To systematically collect essential details, such as author names, paper titles, publication venues, and years, we used a structured data extraction form based on guidelines from Kitchenham et al. (Kitchenham, 2004) and Garousi and Felderer (Garousi and Felderer, 2017). This information was organized in a spreadsheet.
2.5.2. Synthesis and Analysis
We applied thematic analysis to synthesize insights across our research questions, ensuring that coding remained objective and independent of the data extraction structure or biases (Stewart, 2024). Full papers were subsequently imported into ATLAS.ti, a qualitative data analysis tool that allowed for direct text coding. This process adhered to established thematic analysis steps (Braun and Clarke, 2006), providing consistency and depth in our approach. We used a step-by-step approach to systematically identify codes, clusters, and classifications related to our research questions (Kumara et al., 2021):
  • Familiarization with the Data: We began by reading the selected papers to understand the data comprehensively, enabling us to identify the key elements relevant to our research questions.
  • Pilot Study: We randomly selected twenty sources for initial analysis, with the first two authors of the paper independently analyzing each to establish preliminary codes using structural and descriptive coding approaches (Saldaña, 2021). This phase involved creating codes by conceptualizing information related to the research questions. Through iterative review and refinement, researchers adjusted codes to clarify emerging theories and form thematic clusters.
  • Inter-Rater Consensus Analysis for the Pilot Study: After finalizing the coding process, the first two authors of the paper conducted an inter-rater review, assessing each other’s codes to establish consensus on naming, types, and categorization. This collaborative effort helped to refine codes and maintain their consistency.
  • Complete Dataset Coding: The remaining sources were subsequently coded based on the agreement reached during the pilot study. The first author of this paper conducted this step, and the second author reviewed the resulting codes.
  • Inter-Rater Consensus Analysis for the Complete Study: After completing the individual coding, we performed a final inter-rater assessment to ensure the coding aligned with the established framework. Any discrepancies were discussed and resolved.
  • Generating the report: The results of this thematic analysis are presented in Sections 3,  4,  5,  6 and  7.
2.6. Replication Package
For the validation and reproducibility of this work, we have provided an extensive replication package online [^1] https://drive.google.com/drive/folders/1wOu11lB9wsxK7cQKzPL63GAjD9UiiF0s?usp=sharing. This package includes the complete compilation of sources, search queries, and thematic analysis performed using Atlas.ti [^2] https://atlasti.com/ (such as codes, groupings, quotations, and analyses).
3. Chaos Engineering: Definitions, Functionaries, and Elements (RQ1)
This section first examines practitioners’ definitions of chaos engineering. Next, it describes the phases, functional requirements, and non-functional requirements of chaos engineering. Finally, it identifies the main functional components of chaos engineering.
3.1. Definitions
Chaos engineering is a growing practice, but its definition remains inconsistent. Hence, we aim to establish a clear and shared definition by analyzing those proposed by different sources. Originally developed at Netflix, chaos engineering focuses on intentionally introducing faults within a system to assess and enhance its robustness  (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Torkura et al., 2021). However, confusion often arises when distinguishing chaos engineering from traditional testing practices. While conventional testing remains vital, chaos engineering should be seen as a complement rather than a replacement (Green, 2023). Traditional testing and monitoring are crucial, yet they may not be sufficient to expose all possible failure modes in complex systems. Chaos engineering mitigates this issue by proactively identifying the weaknesses of a system that might go unnoticed during standard testing procedures but may surface as failures in production environments (Paul, 2019; Green, 2023).
Table 6. Selected Definitions of Chaos Engineering
Source Definition
Netflix
"the discipline of experimenting on a system to build confidence in its ability to withstand turbulent
conditions in production" (Team, 2024).
Gremlin
"a disciplined approach to identifying failures before they become outages by intentionally
’breaking things’ to build more resilient systems" (Inc., 2023).
AWS
"the process of stressing an application in testing or production environments by creating
disruptive events, such as server outages or API throttling, observing how the system responds,
and implementing improvements" (Starr et al., 2022; Kahil and Arenhage, 2021).
Google "the practice of intentionally introducing failures into production systems under load" (Google Cloud, 2024).
Microsoft "intentionally injecting faults into applications in a controlled manner to simulate real-world issues" (Microsoft, 2023).
IBM
"intentionally causing failures in production or pre-production environments
to understand their impact and improve defense strategies" (IBM, 2024).
Datadog
"a discipline that encourages experimenting in production and injecting controlled failures
into the system to understand how the system will react and improve reliability" (Pulido et al., 2021).
Harness "the science of injecting failure to gauge system resiliency" (Schillerstrom, 2022).
Twilio "intentionally injecting failures to test a system’s ability to self-heal" (Wong, 2017).
Uber "the practice of injecting controlled failures directly into production environments to test resilience" (Jones, 2020).
Splunk
"the practice of assessing the resilience of production systems by intentionally introducing
failures and running experiments" (Wickramasinghe, 2023).
Figure 4. Word Cloud of Key Terms in Chaos Engineering Definitions
Table 6 shows the excerpts from the definitions provided by a subset of the reviewed literature. We also created a word cloud to identify the key terms used in these definitions, as shown in Figure  4. The key themes emphasized in these definitions include system resilience, failure injection, testing in production environments, and controlled experiments. These recurring elements highlight chaos engineering’s focus on deliberately introducing failures in a controlled manner to assess how systems behave under stress and recover from disruptions. Based on these insights, we propose the following unified definition.
Chaos engineering is a resilience testing practice that intentionally injects controlled faults into software systems in production-like or actual production environments to simulate adverse real-world conditions. It enables uncovering hidden vulnerabilities in systems that hinder their ability to withstand potential real-world disruptions. It also enables the assessment and enhancement of systems’ capacity and operational readiness against such disruptions.
In line with the definition, chaos engineering relies on several guiding principles to design and carry out experiments, ensuring that systems are rigorously tested under controlled conditions (Basiri et al., 2016; Aisosa, 2023). Table 7 summarizes these principles.
Table 7. Principles and Measurement Metrics
Principle Description Articles Referenced
Establish Steady State Understand the system’s normal behavior and hypothesize how it will handle chaos (real-world disruptive incidents/failures). (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Gill, 2021; Wickramasinghe, 2023; Cotroneo et al., 2022; Nikolaidis et al., 2021; Frank et al., 2021b; Pierce et al., 2021; Frank et al., 2023; Konstantinou et al., 2021; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Basiri et al., 2019; Zhang et al., 2023; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Bocetta, 2019; Palani and Gupta, 2023; Gill, 2022; Fawcett, 2020; Manna, 2021; Bairyev, 2023; Sbiai et al., 2023; Green, 2023; Krivas and Portela, 2020; Kostic, 2024; Abdul, 2024; Katirtzis, 2022; Le, 2022; Colyer, 2019; Patel, 2022)
Setup Monitoring Infrastructure Set up tools to monitor the system for deviations from its normal behavior during controlled experiments. (Zhang et al., 2019; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Wickramasinghe, 2023; Poltronieri et al., 2022; Zhang et al., 2021a; Nikolaidis et al., 2021; Pierce et al., 2021; Nikolaidis et al., 2023; Malik et al., 2023; Kesim et al., 2020; Siwach et al., 2022; Sbiai et al., 2023; Abdul, 2024; Bairyev, 2023; Green, 2023)
Conduct Chaos Experiments Design and execute experiments to simulate real-world failures. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Gill, 2021; Pierce et al., 2021; Naqvi et al., 2022; Malik et al., 2023; Bedoya et al., 2023; Kesim et al., 2020; Jernberg et al., 2020; Jakkaraju, 2020; Palani and Gupta, 2023; Gill, 2022; Fawcett, 2020; Manna, 2021; Bairyev, 2023; Patel, 2022; Abdul, 2024; Roa, 2022)
Test and Refine Review results, apply improvements, and repeat to boost resilience. (Zhang et al., 2019; Simonsson et al., 2021; Inc., 2023; Bremmers, 2021; Gill, 2021; Wickramasinghe, 2023; Jernberg et al., 2020; Bairyev, 2023; Sbiai et al., 2023; Green, 2023; Kostic, 2024; Colyer, 2019; Roa, 2022; Basiri et al., 2019; Torkura et al., 2020; Fogli et al., 2023; Patel, 2022; Wu, 2021; Tomka, 2024)
Grow Blast Radius Gradually expand experiments to test increasingly complex systems and environments. (Torkura et al., 2020; Zhang et al., 2019; Simonsson et al., 2021; Inc., 2023; Gill, 2021; Wickramasinghe, 2023; Pierce et al., 2021; Frank et al., 2023; Poltronieri et al., 2022; Ikeuchi et al., 2020; Nikolaidis et al., 2021; Naqvi et al., 2022; Malik et al., 2023; Kesim et al., 2020; Zhang et al., 2023; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Fawcett, 2020; Manna, 2021; Bairyev, 2023; Sbiai et al., 2023; Tomka, 2024; Green, 2023; Kostic, 2024)
Expand to Production Run experiments in production to build confidence in the system’s ability to endure adverse real-world events. (Nombela, 2023; Krishnamurthy, 2021; Moon, 2022; Simonsson et al., 2021; Pierce et al., 2021; Konstantinou et al., 2021; Park et al., 2019; Fawcett, 2020)
3.2. Functionalities of Chaos Engineering
Based on our review of academic and grey literature, we identified five key features that define an effective chaos engineering solution: experimentation hub, collaborative workflow management, observability suite, security and governance, and integration and automation (see Figure  5). The rest of this section discusses each in detail.
3.2.1. Experimentation Hub
Experimentation is a crucial aspect of a chaos engineering platform, allowing users to meticulously plan and seamlessly conduct chaos experiments (Katirtzis, 2022; Abdul, 2024; Krivas and Portela, 2020). In this phase, users first define the scope, targets, and parameters of chaos experiments with precision, ensuring a clear focus for testing (Kostic, 2024; Bremmers, 2021; Kadikar, 2023; Wu, 2021; Cotroneo et al., 2022; Katirtzis, 2022; Torkura et al., 2020). The next step is to establish the steady state of the system - baseline performance metrics and normal behavior - which serves as a reference point for evaluating the impact of disruptions (Frank et al., 2023; Konstantinou et al., 2021; Jernberg et al., 2020; Siwach et al., 2022). Based on this steady state, users form hypotheses about how the system will behave under stress, guiding both the design of experiments and the setting of resilience expectations  (Kostic, 2024; Zhang et al., 2019; Wickramasinghe, 2023; Fogli et al., 2023; Dedousis et al., 2023). Once the experiment is defined, the chaos engineering platform injects faults, such as network latency, packet loss, or server failures. This enables users to assess the weaknesses and operational readiness of a system by evaluating its fault handling capabilities (Kostic, 2024; Chen et al., 2024; Fogli et al., 2023; Zhang et al., 2023). As these faults are injected and system behavior is observed, careful planning and execution help chaos experiments effectively identify and address system vulnerabilities, ultimately strengthening overall resilience (Nikolaidis et al., 2021; Colyer, 2019).
3.2.2. Collaborative Workflow Management
Chaos engineering is a collaborative practice that brings together stakeholders such as developers, testers, operations teams, and security experts. These individuals must work closely to design, execute, and analyze experiments to identify system vulnerabilities. Through effective collaborative workflow management, teams can maintain clear communication, coordinate tasks efficiently, and ensure that all parties are informed throughout the experimentation process (Siwach et al., 2022; Dua, 2024; Green, 2023).
Facilitating Team Communication and Coordination.
Chaos engineering relies on continuous communication between relevant parties to ensure alignment with the objectives of the experiment and the expected results (Gill, 2021). By forming cross-functional working groups, participants can discuss scenarios, hypothesize outcomes, and maintain a shared understanding of the objectives of the experiment. This approach ensures that everyone stays coordinated and clear on their roles during the execution and analysis phases. For example, in a simulated database failure, developers can focus on diagnosing and resolving the issue while the operations group works to reroute traffic and maintain service availability. At the same time, security experts assess the potential vulnerabilities that the failure exposes, and testers verify the system response to failures (Patel, 2022). This collaborative process helps avoid information silos and supports collective learning, leading to better preparation for future incidents (Green, 2023; Kadikar, 2023).
Notify Stakeholders about Experiment Progress.
Keeping stakeholders informed about the status of the chaos engineering experiment is crucial to transparency and effective decision making. First, well-designed notification systems trigger alerts based on anomalies, such as unexpected error rates or system performance changes  (Bremmers, 2021). Then, these alerts are automatically sent to stakeholders through integrated notification features. These alerts enable quick issue communication and prompt responses to performance degradation or system instability. In addition, they provide real-time insight into system behaviors, facilitating adjustments during the experiment. Finally, this continuous feedback loop refines the monitoring systems and improves incident preparedness through improved alert accuracy (Shah, 2021).
Maintain Documentation and Knowledge Repositories.
Building a consistent knowledge base for future reference is essential for the long-term success and continuous improvement of chaos engineering practices (Vinisky, 2024; Kadikar, 2023). By maintaining detailed records of experiment scenarios, findings, and recommendations, organizations create a valuable foundation for troubleshooting and informed decision-making (Kadikar, 2023; Gill, 2022). This process should include standardized templates for experiments and comprehensive postmortem reports, ensuring that the documentation is consistent and thorough (Siwach et al., 2022). A well-organized knowledge base preserves essential information and fosters a deeper understanding of system behavior under stress  (Gill, 2021). These insights from previous experiments can directly improve system design and operational practices (Vinisky, 2024). Through consistent documentation and structured knowledge sharing, chaos engineering platforms ensure that lessons learned are preserved, actively shaping future experiments, and strengthening overall system performance (Torkura et al., 2020).
3.2.3. Observability Suite
Chaos engineering platforms use an observability suite to monitor and analyze systems during experiments, allowing teams to detect issues early, identify root causes, and take corrective actions (Green, 2023; National Australia Bank, 2020; Chen et al., 2022). Key features discussed in the following subsections include real-time dashboards and experiment impact analysis to identify faults’ root causes.
Provide Real-Time Dashboards.
Real-time observability depends on dashboards, metrics, traces, and logs for the visualization of experiments. (National Australia Bank, 2020; Gatev, 2021; Pulcinelli et al., 2023). First, dashboards, integrated with monitoring tools, can show critical metrics such as latency, throughput, error rates, and resource usage, allowing engineers to monitor system performance and quickly detect problems (Wickramasinghe, 2023; Manna, 2021; Fawcett, 2020; Green, 2023). Compared to performance metrics, traces can capture the flow of requests through the system, revealing interactions among components and identifying bottlenecks (Chen et al., 2024; Zhang et al., 2021a). Distributed tracing tools allow engineers to analyze component interactions and inefficiencies during chaos experiments (Konstantinou et al., 2021; Frank et al., 2021b). Finally, logs record system events, supporting diagnostic efforts, especially in security contexts (Klinaku et al., 2022; Chen et al., 2022; Green, 2023). Log analysis tools enable engineers to trace anomalies and maintain a complete audit trail  (Torkura et al., 2021; Soldani and Brogi, 2021; Vu et al., 2022).
Analyze Impact and Identify Root Causes of Faults.
After gathering and visualizing data through real-time dashboards, the next crucial step is to analyze the results of chaos experiments and identify the root causes of any faults, using each pillar of observability in the process  (Yu et al., 2021; Gianchandani et al., 2022) The metrics collected during chaos experiments are compared with baseline performance to detect deviations that may indicate vulnerabilities. These deviations help identify how well the system handled stress and where failures occurred (Ji et al., 2023). Logs and trace data provide deeper insight into how individual components interacted and responded to stress. This detailed data helps identify the source of failures and component interactions that contribute to performance problems (Green, 2023). These metrics, logs, and traces can be combined to facilitate root cause analysis and offer qualitative and quantitative insights. For example, this helps engineers determine whether failures were caused by bottlenecks, resource mismanagement, or faulty component interactions (Yu et al., 2021). By understanding the root causes of failures, teams can implement targeted improvements to enhance and prevent future failures (Gill, 2022; Nombela, 2023; Gianchandani et al., 2022).
3.2.4. Security and Governance
Chaos engineering platforms require a robust security and governance framework to ensure that experiments are conducted safely and comply with regulatory standards (Fogli et al., 2023; Kostic, 2024). This section outlines vital mechanisms that protect system integrity while adhering to industry regulations, focusing on implementing safety controls, automated rollbacks, and ensuring regulatory compliance.
Implement Safety Controls and Automated Rollbacks.
Our literature identifies several vital mechanisms, including abort conditions, automated rollbacks, and blast radius control, which ensure that experiments are conducted safely (Patel, 2022).
  • Abort conditions: These ensure that chaos experiments are immediately stopped if predefined thresholds are exceeded (Abdul, 2024; Sbiai et al., 2023). For example, the experiment is automatically terminated if performance metrics, such as latency or availability, degrade beyond acceptable levels. This prevents further damage and maintains system stability by stopping experiments before they can cause widespread problems (Torkura et al., 2020).
  • Automated rollbacks: These allow systems to revert to a known stable state after an experiment (Fawcett, 2020; Miles, 2019b). In cases where a simulated failure causes service disruption, automated rollback mechanisms quickly restore the system to its pre-experiment configuration. This reduces downtime and prevents prolonged instability in the system (Gill, 2022).
  • Blast radius control: This mechanism limits the scope of chaos experiments to isolated parts of the system (Wickramasinghe, 2023; Manna, 2021). For example, teams might start by targeting a single microservice instead of testing the entire infrastructure. Gradually expanding the blast radius allows controlled testing, ensuring vulnerabilities can be identified without risking widespread system failures (Abdul, 2024). Additionally, Gremlin  (Inc., 2023) recommends using the Rumsfeld matrix, a framework to categorize system behaviors during chaos experiments. This approach enables teams to understand better and anticipate various types of system responses by classifying them as follows:
    • Known Knowns: Aspects of the system that are fully understood and predictable.
    • Known Unknowns: Recognized uncertainties that pose potential risks.
    • Unknown Knowns: Understood risks that may not be immediately apparent or considered.
    • Unknown Unknowns: Completely unforeseen issues that may emerge during experimentation.
Starting with known knowns, teams establish a baseline of predictable system behavior. They can then progressively move into more uncertain areas like Known Unknowns and Unknowns Knowns before finally addressing the most unpredictable Unknown Unknowns (Palani and Gupta, 2023). This step-by-step approach reduces system risks by building understanding before testing less predictable areas.
Ensure Regulatory Compliance.
Adhering to industry standards and legal frameworks is essential to ensure that systems meet established data protection and security requirements, including guidelines such as the General Data Protection Regulation (GDPR) and IEC 62443 (Helmke, 2019; Torkura et al., 2021). These frameworks provide organizations with critical guidance to safeguard sensitive information, maintain operational integrity, and protect vital infrastructures (Helmke, 2019; Kostic, 2024). To meet these compliance requirements, some organizations adopt Infrastructure as Code (IaC), which automates the provisioning and management of infrastructure (Torkura et al., 2020; Fogli et al., 2023). Tools such as Terraform and Ansible are commonly used in IaC to ensure consistent implementation of security policies, minimize human error, and maintain infrastructure in alignment with regulatory standards (Torkura et al., 2021; Vu et al., 2022). In addition to IaC, monitoring tools such as Datadog  (Abdul, 2024; Mooney, 2023), which detects compliance violations and configuration drift such as publicly exposed data, and Security Monkey, which monitors Secure Sockets Layer (SSL) certificates and other security settings to prevent vulnerabilities such as certificate expiration, are essential to maintain compliance (Dua, 2024; Tivey et al., 2019).
3.2.5. Integration and Automation with Continuous Integration and Continuous Delivery (CI/CD)
CI/CD pipelines are critical to optimizing the efficiency and consistency of chaos experiments (Rivera et al., 2023; Bairyev, 2023; Azure Readiness, 2023). When performed manually, tasks such as performing experiments and monitoring system behavior can be inefficient and prone to human error (Kostic, 2024; Jernberg et al., 2020; Bremmers, 2021; Chen et al., 2024). Automating these processes minimizes human intervention, ensuring tasks are performed reliably and consistently, thus reducing the likelihood of mistakes (Wickramasinghe, 2023; Camacho et al., 2022; Simonsson et al., 2021; Green, 2023). Integrating chaos experiments into CI/CD pipelines further enhances this reliability by automatically initiating tests as part of the routine development workflow (Roa, 2022). With each new deployment or configuration update, chaos tests run alongside other routine tests, enabling early detection of vulnerabilities. This proactive approach helps ensure issues are addressed in the development phase, resulting in a more comprehensive and practical testing process (Bremmers, 2021; Manna, 2021)
Figure 5. Key Functionalities Provided by Chaos Engineering.
[…]
Functional Requirement
3.3. Quality Requirements for Chaos Engineering Platform
We have identified 12 quality requirements for a chaos engineering platform from the selected literature: performance, recoverability, resilience, timeliness, observability, scalability, effectiveness, robustness, accuracy, fault tolerance, cost, and adaptability. We provide these requirements and the corresponding metrics for measuring them in Table 8.
Table 8. Quality Attributes and Measurement Metrics
Quality Attribute Measurement Metrics Articles Referenced Total
Performance Evaluate resource use and performance under attack, ensuring chaos experiments align with production metrics. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Nikolaidis et al., 2021; Vu et al., 2022; Klinaku et al., 2022) 10
Recoverability Measure mean time to repair (MTTR), restoration speed, and detection response. (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Patel, 2022; Gill, 2021; Cotroneo et al., 2022; Camacho et al., 2022; Zhang et al., 2021a; Nikolaidis et al., 2021; Pierce et al., 2021; Vu et al., 2022; Frank et al., 2021b; Malik et al., 2023; Zhang et al., 2023; Tomka, 2024; Kadikar, 2023) 16
Resilience Tolerance for degraded performance under stress; success rate of automated restoration actions. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Inc., 2023; Bremmers, 2021; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Vu et al., 2022; Nikolaidis et al., 2023; Frank et al., 2023; Kassab et al., 2022; Konstantinou et al., 2021; Nombela, 2023; Krishnamurthy, 2021; Bocetta, 2019; Manna, 2021; Kostic, 2024; Krivas and Portela, 2020; National Australia Bank, 2020; Azure Readiness, 2023) 20
Timeliness Time to restore service and detect interruptions; impact on user experience and disruption. (Camacho et al., 2022; Pierce et al., 2021; Vu et al., 2022; Frank et al., 2021b; Malik et al., 2023; Cotroneo et al., 2022; Basiri et al., 2019; Gill, 2021; Zhang et al., 2019; Durai, 2022; Shah, 2021; Katirtzis, 2022; Dua, 2024; Wu, 2021; Palani and Gupta, 2023; Manna, 2021) 16
Observability Assess logging completeness and monitoring uptime; measure alert latency and tracing success rate. (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Camacho et al., 2022; Torkura et al., 2021; Zhang et al., 2021a; Vu et al., 2022; Frank et al., 2023; Malik et al., 2023; Kesim et al., 2020; Siwach et al., 2022; Konstantinou et al., 2021; Hawkins, 2020; Manna, 2021; Kostic, 2024; Kadikar, 2023; Azure Readiness, 2023) 20
Scalability Manage increasing attack rates and scale experiment complexity. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Patel, 2022; Gill, 2021; Vu et al., 2022; Sondhi et al., 2021; Malik et al., 2023; Klinaku et al., 2022; Camilli et al., 2022; Nikolaidis et al., 2023; Kassab et al., 2022; Rivera et al., 2023; Krishnamurthy, 2021; Durai, 2022; Wu, 2021; Palani and Gupta, 2023; Manna, 2021; Azure Readiness, 2023; Vinisky, 2024; Kostic, 2024) 20
Effectiveness Measure achievement rates, end-user satisfaction, and time to achieve outcomes. (Sondhi et al., 2021; Chen et al., 2022; Malik et al., 2023; Wu, 2021) 4
Robustness Evaluate stability at peak load and error tolerance. (Vu et al., 2022; Ma et al., 2023; Yu et al., 2021; Wu, 2021; Bedoya et al., 2023; Chen et al., 2022; Vinisky, 2024) 7
Accuracy Evaluate system consistency through recovery accuracy and fault localization precision. (Green, 2023; Simonsson et al., 2021; Chen et al., 2022) 3
Fault Tolerance Automated failover success rate during faults; immediate recovery time after detected faults. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Gill, 2021; Wickramasinghe, 2023; Zhang et al., 2021a; Camacho et al., 2022; Poltronieri et al., 2022; Nikolaidis et al., 2021; Ma et al., 2023; Meiklejohn et al., 2021; Sondhi et al., 2021; Azure Readiness, 2023; Tomka, 2024; Manna, 2021) 15
Cost Cost of chaos engineering and savings from preventing outages. (Gill, 2021; Patel, 2022; Kesim et al., 2020; Palani and Gupta, 2023; Klinaku et al., 2022; Inc., 2023; Bocetta, 2019; Basiri et al., 2019; National Australia Bank, 2020) 9
Adaptability Change methods and platforms for new systems. (Vu et al., 2022; Ma et al., 2023; Yu et al., 2021; Wu, 2021) 4
3.4. Key Components of Chaos Engineering
Before adopting a chaos engineering framework, organizations should carefully evaluate its key components to ensure that they align with their operational requirements and objectives. We have identified several essential components for an effective chaos engineering platform. Most of the reviewed studies incorporate a combination of these components, which we have categorized into five central units: experimentation, collaboration and communication, observability and analysis, security and governance, and automation and integration with CI/CD. This classification, derived from the functionalities discussed in Section  3.2, is shown in Figure  6. The remainder of this section provides a detailed analysis of each component within this framework.
Figure 6. Key Components of Chaos Engineering.
[…]
Key Components of Chaos Engineering.
3.5. Experimentation Unit
Successful coordination of chaos experiments depends on a systematic approach to both planning and execution (Zhang et al., 2021a; Nikolaidis et al., 2021). To this end, the experimentation unit is built around two foundational components: the experiment designer and the failure injection hub, which will be explored further in this section.
3.5.1. Experiment Designer
This module is crucial for structuring chaos experiments, allowing teams to define which systems to target (Zhang et al., 2021a; Camacho et al., 2022). After identifying the systems, the teams can choose the fault types most relevant to their testing scenarios, such as network failures or system crashes (Zhang et al., 2021a). Once the faults are defined, the experiment designer helps organize the timing and order in which they are introduced, ensuring that they reflect real-world conditions (Camacho et al., 2022). By enabling these customized failure simulations, the experiment designer ensures that the tests are well-structured, providing consistent data for postmortem analysis (Zhang et al., 2021a; Konstantinou et al., 2021). This structure allows teams to track how the behavior of the system evolves under different failure conditions. For example, teams might simulate a server crash during peak traffic to test failover capabilities and measure the impact on user experience (Zhang et al., 2019; Simonsson et al., 2021).
3.5.2. Fault Injection Hub
Once the experiment is designed and the faults are organized, the next phase is to inject these controlled faults and workloads into the system, simulating real-world scenarios (Fawcett, 2020; Palani and Gupta, 2023). The fault injection process, as outlined by Hseh et al (Hsueh et al., 1997), in Figure  7, involves multiple components working together: a controller, fault injector, fault library, workload generator, workload library, and target system (Nikolaidis et al., 2021; Frank et al., 2021b). Each of these components plays a critical role in ensuring the accuracy and smooth execution of the chaos experiment. Central to this process is the controller, which manages the introduction of faults and dynamically adjusts based on real-time system feedback, ensuring that the experiment remains aligned with its objectives (Torkura et al., 2020). As part of the hub, the fault injector introduces specific disruptions, such as hardware or software failures, selected from a diverse fault library that offers teams a variety of failure scenarios to choose from (Soldani and Brogi, 2021; Frank et al., 2021b). In parallel, the workload generator creates synthetic workloads based on predefined patterns stored in the workload library. This allows teams to simulate normal or peak traffic conditions to observe how the system handles varying stress levels (Fogli et al., 2023; Kesim et al., 2020). These faults and workloads are then applied to the target system, be it a production or preproduction environment, providing teams with valuable information on the weaknesses of the system and the recovery capabilities (Wickramasinghe, 2023; Zhang et al., 2021a).
Figure 7. Conceptual Schema of Fault Injection(Hsueh et al., 1997).
[…]
Conceptual schema of fault injection.
3.6. Collaboration and Communication Unit
Executing chaos experiments across teams requires strong communication and access to data from previous experiments (Kadikar, 2023; Patel, 2022; Siwach et al., 2022), which are supported by the alert and notification system and the knowledge base.
3.6.1. Alert and Notification System
The hub integrates chaos tools with existing alert systems to ensure smooth communication during chaos experiments. This integration ensures that teams are immediately notified of critical events such as experiment initiation, fault injection, and any unexpected behaviors during the process (Starr et al., 2022; Vinisky, 2024). These real-time notifications are distributed through multiple channels, including email and Slack, ensuring that all stakeholders receive timely updates. For example, if an issue arises, the system immediately alerts the relevant individuals, allowing them to respond quickly and address the problem effectively (Torkura et al., 2020; Shah, 2021). This prompt response mechanism is critical for maintaining system stability, as it helps mitigate potential risks before they escalate (Siwach et al., 2022).
3.6.2. Knowledge Base
Maintaining a well-organized archive of previous chaos experiments is crucial to continuously improve the resilience of the system and support real-time communication between users (Kadikar, 2023). This knowledge base is a central repository that stores comprehensive documentation that includes experiment designs, execution procedures, and results (Zhang et al., 2019; Torkura et al., 2020). Recalling past experiments, teams can leverage valuable insights from successful and unsuccessful tests, fostering a cycle of learning and enhancement. Detailed post-mortem reports, for example, allow teams to thoroughly analyze the results, identify improvement areas, and strengthen the resilience of the system (Inc., 2023; Palani and Gupta, 2023). Furthermore, using standardized templates ensures that this knowledge is systematically preserved, promoting consistency and reliability in future chaos experiments (Siwach et al., 2022; Zhang et al., 2019).
3.7. Observability and Analysis Unit
Monitoring system behavior during chaos experiments is crucial for understanding faults’ impact and identifying areas for improvement (Gianchandani et al., 2022; Nikolaidis et al., 2021; Yu et al., 2021). Essential components, such as data collector and data analyzer, play a vital role in this process, ensuring that the systems are evaluated and optimized effectively (Zhang et al., 2021a; Simonsson et al., 2021).
3.7.1. Data Collector
Effective data collection is essential during chaos experiments, capturing various types of information such as performance metrics, logs, and traces (Gatev, 2021). These data points clearly explain how the system responds under stress (Zhang et al., 2021a). The data collector gathers performance metrics to track key indicators such as response times, error rates, latency, and system throughput. These data are crucial to measuring the effectiveness of the system in detecting faults  (Cotroneo et al., 2022). Logs document significant events within the system, including errors and state transitions. These records help diagnose issues and improve the understanding of system behavior during failures  (Vu et al., 2022; Cotroneo et al., 2022). Tracing methods can monitor the flow of requests across distributed services, enabling the identification of bottlenecks and observing how traffic is managed under stress (Zhang et al., 2021a). All data collected are visualized through a centralized dashboard, providing a straightforward and organized interface that allows quick assessment of performance indicators and overall system health (Malik et al., 2023; Basiri et al., 2019; Camacho et al., 2022).
3.7.2. Data Analyzer
The information collected during chaos experiments is systematically analyzed to identify essential trends and patterns (Zhang et al., 2021a; Simonsson et al., 2021). This process involves impact analysis, which evaluates how failures influence system performance based on previously collected metrics (Pulcinelli et al., 2023). By analyzing these data, it is possible to determine how disruptions affect the system and identify its most vulnerable components  (Fawcett, 2020; Cotroneo et al., 2022). Additionally, root cause analysis leverages logs and traces to trace failures back to their origins, whether due to a configuration error, service malfunction, or hardware fault (Zhang et al., 2021a; Cotroneo et al., 2022).
3.8. Security and Governance Unit
Ensuring the secure execution of chaotic experiments in production environments is essential to prevent disruptions and unauthorized actions, achieved through adequate safety and access control mechanisms (Al-Said Ahmad et al., 2024; Green, 2023; Gianchandani et al., 2022). First, safety controls mitigate the risk of chaos experiments that cause severe outages by incorporating predefined thresholds, such as automated rollbacks, to restore systems to a steady state if an experiment causes prolonged disruptions (Zhang et al., 2021a; Gill, 2022). These controls allow the system to automatically abort experiments when severe impacts are detected, maintaining overall system stability (Kadikar, 2023). By limiting the blast radius to specific areas, safety controls ensure that experiments do not cause widespread failures (Simonsson et al., 2021; Gill, 2021), with fail-safe conditions ensuring immediate containment of issues (Bhaskar, 2022). Second, access control mechanisms further enhance security by restricting the ability to initiate, modify, or stop chaos experiments only to authorized personnel (Torkura et al., 2020; Sbiai et al., 2023). This ensures that only individuals with the appropriate permissions can execute or alter experiments, particularly in sensitive environments where unauthorized changes could lead to significant security and compliance risks  (Fogli et al., 2023; Tivey et al., 2019).
3.9. Automation and Integration Unit
This unit is vital in embedding chaos engineering into the software development lifecycle. It automates and integrates chaos experiments with existing systems to ensure consistent and efficient execution. First, the unit connects the chaos engineering platform with essential tools such as CI/CD pipelines, cloud platforms, and monitoring systems. This integration allows chaos experiments to be seamlessly incorporated into routine workflows such as deployments, updates, and system monitoring, enabling automatic execution as part of regular operations (Zhang et al., 2021a; Torkura et al., 2020). Second, the unit automates the execution of chaos experiments and includes a scheduling mechanism that allows tests to run at predefined intervals, such as during off-peak hours or after system updates (Cotroneo et al., 2022; Zhang et al., 2019). This ensures that chaos experiments are performed regularly and reliably, reducing the risk of human error and allowing organizations to test their systems without interrupting normal activities (Camacho et al., 2022; Simonsson et al., 2021).
4. Motivation Behind Chaos Engineering (RQ2)
This section delves into the challenges driving chaos engineering adoption in organizations. Figure 8 shows the technical and socio-technical challenges we identified from the reviewed literature.
Figure 8. Challenges that Drive Chaos Engineering Adoption.
[…]
Challenges that drive chaos engineering adoption.
4.1. Technical Challenges
Chaos engineering aims to tackle various technical challenges that undermine system resilience. The following subsections explore these challenges in detail.
Unanticipated System Failures
In a distributed system, one of the most critical risks is the occurrence of unexpected failures in various parts of the system (Ahmed and Wu, 2013). Given the highly interconnected nature of modern systems, a failure in one component can quickly cascade into widespread disruptions across multiple services or functions (Wickramasinghe, 2023). However, traditional testing methods often struggle to identify deep-rooted system vulnerabilities because they focus on known scenarios and isolated components, neglecting real-world systems’ dynamic and complex interactions (Soldani and Brogi, 2021). In contrast, chaos engineering adopts a proactive approach by deliberately introducing failures, such as server crashes, network interruptions, and resource exhaustion, in environments that closely resemble production. As a result, engineers can observe how the system behaves under stress, thus revealing hidden weaknesses that conventional testing methods might overlook. By focusing on unpredictable interactions between distributed components, chaos engineering effectively uncovers vulnerabilities that arise from the complexity of the system (Bairyev, 2023). For example, chaos experiments might simulate network partitioning between microservices, which can reveal issues related to data consistency and service availability (Bairyev, 2023; Fogli et al., 2023). These controlled disruptions allow engineers to better understand how different system components interact under turbulent conditions (Torkura et al., 2020).
Inadequate Fault Tolerance Mechanisms
Many distributed systems are built with fault tolerance mechanisms to handle routine issues, such as server failures or network delays (Inc., 2023). However, when larger-scale failures occur, e.g., multi-region outages or prolonged network disruptions, these mechanisms often fail to maintain system stability (Palani and Gupta, 2023; Silveira, 2023; Nallainathan, 2023). This shortfall occurs because mechanisms designed to handle more minor issues may not be able to handle the broader impact of significant failures, leading to severe performance degradation or even total system outages (Basiri et al., 2019; Gill, 2021). Chaos engineering addresses these shortcomings by targeting and testing fault tolerance mechanisms in scenarios replicating the most significant risks to system stability (Tivey et al., 2019). For example, chaos engineering can simulate the unavailability of critical services in one region to determine whether traffic is adequately redirected to the other areas or if the load balancers of the system can effectively distribute incoming requests between redundant resources (Bremmers, 2021; Aharon, 2024). Furthermore, by testing the ability of backup systems to maintain service levels during prolonged outages, chaos engineering reveals whether these fallback mechanisms are sufficient for extended failures (Inc., 2023; Wickramasinghe, 2023; Basiri et al., 2019). These focused experiments help ensure systems can handle severe disruptions without experiencing widespread failure (Bairyev, 2023).
Limited Performance Insights and Failure Assessment
Another significant challenge in distributed systems is understanding how they perform during high-stress or failure scenarios (Chen et al., 2006; Fidge, 1996). When systems operate under normal conditions, traffic bottlenecks, inefficient resource usage, or service communication failures often remain hidden (Fogli et al., 2023). This lack of visibility leaves teams unprepared for unexpected disruptions, which can lead to outages or performance degradation when systems are stressed (Konstantinou et al., 2021). Chaos engineering addresses this by introducing specific disruptions, such as intentionally increasing network latency or restricting memory resources on critical components (Siwach et al., 2022). These controlled failures are monitored through observability tools that provide real-time visibility into metrics such as latency, error rates, and resource consumption (Wachtel, 2023). By analyzing these metrics, teams gain detailed insight into how services handle stress - uncovering issues such as delayed responses or inefficient resource utilization that would otherwise remain hidden during normal operations (Bremmers, 2021; Palani and Gupta, 2023).
4.2. Socio-Technical Challenges
Chaos engineering can address various socio-technical challenges, expanding its impact beyond technical issues to encompass the complex interplay between people, processes, and technology (Miles, 2019b; Thomas et al., 2019; Patriarca et al., 2018).
Inconsistent Experiment Design and Execution
When experiment design and execution lack consistency, teams struggle to accurately assess the behavior of the system under stress (Farias et al., 2012; Caferra et al., 2023). Factors such as varying test timing, failure scenarios, or changing system states lead to unreliable data. This makes it difficult to identify critical vulnerabilities or adequately assess the system’s ability to handle disruptions (Klinaku et al., 2022). Without standardization, essential insights into system responses may be overlooked, creating false confidence in its resilience (Danilowicz and Nguyen, 2003). Chaos engineering addresses this by introducing a structured and repeatable approach, ensuring that predefined failure scenarios such as network issues or resource depletion are applied consistently across tests (Zhang et al., 2021a; Naqvi et al., 2022; Kostic, 2024; Vu et al., 2022). This uniformity eliminates variability, producing reliable, comparable data that allows teams to pinpoint weaknesses and confidently evaluate recovery mechanisms (Kadikar, 2023; Azure Readiness, 2023).
Lack of Collaboration and Knowledge Sharing
Limited collaboration and challenges in knowledge sharing between teams, such as development, operations, and security, can lead to communication gaps and the formation of silos, where vital information may not be effectively exchanged (Habeh et al., 2021; Riege, 2005). This results in misaligned responses and repeated errors during system failures, which ultimately slows recovery efforts and misses opportunities for improvement (de Vreede et al., 2016). Chaos engineering tackles this issue by fostering cross-functional collaboration, engaging all relevant teams in both the planning and execution of chaos experiments (Tomka, 2024; Tivey et al., 2019; Azure Readiness, 2023). For example, when simulating a specific failure, such as a service outage or database failure, development teams define the expected behavior of the system, operations teams monitor the system’s performance under stress, and security teams evaluate potential vulnerabilities exposed during the failure (Krishnamurthy, 2021). This multi-team involvement ensures that the insights from the experiment, whether an unexpected failure response or a missed alert, are immediately shared and understood by everyone (Green, 2023; Katirtzis, 2022; Dua, 2024). Furthermore, the results of each chaos experiment are documented and discussed in collaborative postmortem sessions, where all teams contribute to analyzing what went wrong, how to fix it, and how to prevent similar issues in the future (Siwach et al., 2022).
Lack of Risk Management
Effective risk management is critical in both production and non-production environments, yet many systems are exposed to unexpected failures due to insufficient preparation (Bremmers, 2021; Naqvi et al., 2022). In non-production settings, such as development or staging, failure scenarios are often inadequately tested, leading to a false sense of security (Wickramasinghe, 2023; Jernberg et al., 2020). These environments typically lack the complexity of live production systems, meaning that when software is deployed, it may not be fully equipped to handle adverse real-world events (Green, 2023). Consequently, when systems reach production, where they face real user traffic, external dependencies, and unpredictable conditions, even minor issues can escalate to significant outages, resulting in costly downtime and degraded user experiences (Gill, 2021; Zhang et al., 2019; Ji et al., 2023). Chaos engineering solves these challenges by enabling controlled experimentation that tests system behavior under simulated and real-world failure conditions (Torkura et al., 2020; Kostic, 2024). In production environments, controlled blast radiuses ensure that failures are confined to isolated parts of the system, thereby preventing widespread disruption (Palani and Gupta, 2023; Siwach et al., 2022; Gianchandani et al., 2022; Mooney, 2023; Bocetta, 2019). Predefined abort conditions and rollback procedures allow teams to stop experiments and recover quickly if the stability of the system is threatened (Abdul, 2024; Sbiai et al., 2023; Siwach et al., 2022).
Incident Management and Improvement
In complex, distributed systems, incident management faces significant challenges due to the unpredictability of failures and the limitations of reactive strategies (Gianchandani et al., 2022). In these approaches, teams respond to incidents only after they occur, focusing on diagnosing and resolving issues after disruptions have already impacted operations (Bremmers, 2021). This often leads to prolonged downtime and repeated failures, as reactive methods limit the opportunities to address these problems proactively (Patel, 2022). Chaos engineering offers a proactive approach by simulating targeted failures that stress critical parts of the system, revealing weaknesses in the incident response process (Anwar and Arunachalam, 2019; Siwach et al., 2022). For example, simulating a sudden network partition might show whether monitoring systems detect the issue promptly and if the system can reroute traffic as intended (Vinisky, 2024; Green, 2023; Gill, 2021). Similarly, inducing a Central Processing Unit (CPU) or memory exhaustion event could expose delays in triggering alerts or inefficiencies in the automated recovery process (Camacho et al., 2022; Kadikar, 2023; Palani and Gupta, 2023; Roa, 2022). By creating these controlled failure scenarios, chaos engineering helps teams identify where their incident response protocols fall short, whether in monitoring, alert escalation, or failover processes (Wu, 2021; Zhang et al., 2021a). This enables teams to proactively adjust their strategies, improving their ability to handle actual incidents faster, more accurately, and with minimal disruption (Dua, 2024; Rivera et al., 2023).
4.3. Benefits
Table 9 outlines the organizational benefits derived from the functionalities described in Section 3.2, addressing the previously identified challenges. These functions include defining chaos experiments, facilitating team communication, notifying stakeholders, maintaining documentation, and using real-time monitoring dashboards. The corresponding activities, such as establishing experimental parameters, analyzing system performance metrics, and fostering cross-functional collaboration, contribute to several organizational advantages. These include improved system understanding, enhanced error-handling capabilities, vital team coordination during stress conditions, and proactive system tuning. Furthermore, automation of tasks and real-time system monitoring can minimize disruptions and optimize system performance and user experience.
Table 9. Mapping of Key Activities Performed by Chaos Engineering with Organizational Benefits
Key Functions Activity Performed Benefits of Organization Articles Total
Define Chaos Experiments Set experiment scope, define normal conditions, hypothesize, and inject fault. Gain system insights, refine the design, establish a baseline, and test error handling. (Kostic, 2024; Gill, 2021; Manna, 2021; Inc., 2023; Wickramasinghe, 2023; Mooney, 2023; Fawcett, 2020; Vinisky, 2024; Al-Said Ahmad et al., 2024; Pierce et al., 2021; Torkura et al., 2020; Vu et al., 2022; Bedoya et al., 2023; Siwach et al., 2022; Zhang et al., 2019; Le, 2022; Palani and Gupta, 2023; Soldani and Brogi, 2021; Azure Readiness, 2023; Dedousis et al., 2023; Colyer, 2019; Patel, 2022; Shah, 2021; Green, 2023; Fogli et al., 2023; National Australia Bank, 2020; Cotroneo et al., 2022; Konstantinou et al., 2021; Zhang et al., 2023; Bairyev, 2023) 30
Facilitate Team Coordination Create groups for team collaboration. Strengthens coordination and insight sharing. (Gill, 2021; Manna, 2021; Dua, 2024; Kostic, 2024; Katirtzis, 2022; Krishnamurthy, 2021; Torkura et al., 2020; Green, 2023; Frank et al., 2021b; Kesim et al., 2020; Bocetta, 2019; Bremmers, 2021) 12
Notify Stakeholders Update stakeholders on progress and results. Promote transparency and prompt updates. (Gill, 2021; Manna, 2021; Dua, 2024; Kostic, 2024; Katirtzis, 2022; Krishnamurthy, 2021; Torkura et al., 2020; Green, 2023; Frank et al., 2021b; Kesim et al., 2020; Bocetta, 2019; Bremmers, 2021) 12
Maintain Knowledge Repositories Share findings for organizational learning. Create a knowledge base to analyze system stress behavior. (Gill, 2021; Manna, 2021; Dua, 2024; Kostic, 2024; Katirtzis, 2022; Krishnamurthy, 2021; Torkura et al., 2020; Green, 2023; Frank et al., 2021b; Kesim et al., 2020; Bocetta, 2019; Starr et al., 2022; Vinisky, 2024; Anwar and Arunachalam, 2019; Kadikar, 2023; Patel, 2022; Vu et al., 2022; Zhang et al., 2019, 2023; Rivera et al., 2023; Bremmers, 2021) 21
Provide Real-Time Dashboards for Visualization Track faults and visualize performance under stress. Assess system health for rapid anomaly detection and refinement. (Tivey et al., 2019; Kadikar, 2023; Pierce et al., 2021; Zhang et al., 2021a; Wickramasinghe, 2023; Gill, 2022; National Australia Bank, 2020; Simonsson et al., 2021; Al-Said Ahmad et al., 2024; Torkura et al., 2020; Gill, 2021; Kostic, 2024; Fogli et al., 2023; Bedoya et al., 2023; Manna, 2021; Krivas and Portela, 2020; Poltronieri et al., 2022; Camacho et al., 2022; Torkura et al., 2021; Pulcinelli et al., 2023; Chen et al., 2022; Yu et al., 2021; Gianchandani et al., 2022; Palani and Gupta, 2023; Anwar and Arunachalam, 2019; Krishnamurthy, 2021; Roa, 2022; Vinisky, 2024; Siwach et al., 2022; Green, 2023; Hawkins, 2020; Vu et al., 2022; Ma et al., 2023; Rivera et al., 2023) 34
Analyze Impact and Identify Root Causes of Faults Assess outcomes to measure impact and analyze root causes. Enhance resilience, identify weaknesses, and improve stability. (Bairyev, 2023; Siwach et al., 2022; Fogli et al., 2023; Zhang et al., 2019; Inc., 2023; Chen et al., 2024; Bhaskar, 2022; Al-Said Ahmad et al., 2024; Gill, 2022; Wickramasinghe, 2023; Yu et al., 2021; Ji et al., 2023; Gianchandani et al., 2022; Nikolaidis et al., 2021; Rivera et al., 2023; Konstantinou et al., 2021; Torkura et al., 2020) 17
Implement Safety Controls and Automated Rollbacks Set up safety mechanisms, automated rollbacks, and abort thresholds to prevent severe disruptions. Minimizes risk, ensures safe experimentation, and reduces downtime with quick recovery measures. (Wickramasinghe, 2023; Siwach et al., 2022; Bocetta, 2019; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Bairyev, 2023; Malik et al., 2023; Inc., 2023; Ikeuchi et al., 2020; Ma et al., 2023; Nikolaidis et al., 2023; Bhaskar, 2022; Simonsson et al., 2021; Gill, 2021; Krishnamurthy, 2021; Torkura et al., 2020; Katirtzis, 2022; Bedoya et al., 2023; Kadikar, 2023; Patel, 2022; Soldani and Brogi, 2021; Frank et al., 2021b; Rivera et al., 2023; Chen et al., 2024; Zhang et al., 2021a, 2019; Chen et al., 2022; Basiri et al., 2019; Bremmers, 2021; Jernberg et al., 2020; Kostic, 2024) 31
Ensure Regulatory Compliance Align chaos experiments with regulatory standards and monitor compliance. Reduce non-compliance risk and build trust by adhering to industry standards. (Wickramasinghe, 2023; Siwach et al., 2022; Bocetta, 2019; Gill, 2022; Fawcett, 2020; Manna, 2021; Gianchandani et al., 2022; Bairyev, 2023; Malik et al., 2023; Inc., 2023; Ikeuchi et al., 2020; Ma et al., 2023; Nikolaidis et al., 2023; Bhaskar, 2022; Simonsson et al., 2021; Gill, 2021; Jernberg et al., 2020; Kadikar, 2023; Kesim et al., 2020; Zhang et al., 2019; Krivas and Portela, 2020; Sbiai et al., 2023; Mooney, 2023; Dua, 2024; Torkura et al., 2021; Palani and Gupta, 2023; Pulcinelli et al., 2023; Kostic, 2024) 28
Integrate with CI/CD and Third-Party Tools Embed chaos experiments in CI/CD with monitoring integration. Automate resilience testing in the software lifecycle to boost efficiency. (Manna, 2021; Azure Readiness, 2023; Konstantinou et al., 2021; Palani and Gupta, 2023; Kostic, 2024; Torkura et al., 2021) 6
Automate Repetitive and Manual Tasks Automate setup and cleanup of chaos experiments using scripts and tools. Reduce time and errors by automating repetitive tasks for consistency and efficiency. (Camacho et al., 2022; Kadikar, 2023; Miles, 2019a; Green, 2023; Nikolaidis et al., 2021; Poltronieri et al., 2022; Fogli et al., 2023; Torkura et al., 2020; Gill, 2021; Cotroneo et al., 2022; Patel, 2022; Bairyev, 2023; Colyer, 2019; Manna, 2021; Zhang et al., 2019; Kostic, 2024; Fawcett, 2020; Jakkaraju, 2020; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Mooney, 2023; Kesim et al., 2020; Zhang et al., 2023; Jernberg et al., 2020; Bremmers, 2021; Torkura et al., 2021; Vinisky, 2024; Chen et al., 2022; Vu et al., 2022) 29
5. Chaos Engineering Framework: Taxonomy, Tools, Practices, and Evaluation (RQ3)
This section presents a taxonomy for chaos engineering and compares existing tools using the taxonomy dimensions. We also present the best practices for adopting chaos engineering in organizations.
5.1. Taxonomy of Chaos Engineering
Figure 9 shows the proposed taxonomy, which organizes the capabilities of a chaos engineering platform across several key dimensions: execution environment, automation strategy, automation mode, deployment type, and evaluation approach. This taxonomy provides organizations with a structured approach to adopting chaos engineering practices by offering clear guidelines for selecting appropriate execution environments, choosing automation strategies that integrate essential tools such as monitoring, logging, and observability, and determining the optimal deployment types (e.g., production or pre-production) based on their risk tolerance. It also helps define task modes (manual, semi-automated, or fully automated). It guides the selection of fault injection tools, ensuring that organizations have the necessary evaluation methods to measure the impact of experiments and derive actionable insights for system improvement.
5.1.1. Execution Environment
This refers to the underlying infrastructure and system components where chaos engineering experiments are performed. These components are categorized into endpoints, grouped by resource types and cloud environments, representing various deployment models. This section will discuss both of these in detail.
Endpoints
These are physical or virtual devices connected to a computer network to exchange data. These include virtual machines (VMs), containers, and serverless functions (Microsoft, 2024). Each type of endpoint presents unique challenges that chaos engineering addresses by simulating failure conditions to improve the robustness of the system and ensure continuous operation.
Virtual machines (VMs) offer a layer of abstraction over physical hardware, allowing multiple independent operating systems to share resources efficiently (Sharma et al., 2016; Smith and Nair, 2005). However, VMs can experience performance issues when resources are overutilized, and they are vulnerable to hypervisor failures and migration difficulties (Azure Readiness, 2023; Fogli et al., 2023). Chaos engineering addresses these challenges by testing VMs under stress conditions such as CPU or memory exhaustion to assess if VMs can recover from failures or seamlessly migrate workloads (Kostic, 2024; Chen et al., 2024; Green, 2023).
Containers, which encapsulate applications and their dependencies in lightweight environments, provide greater efficiency and portability compared to VMs (Sharma et al., 2016; Koskinen et al., 2019). However, containers face resource contention, network instability, and orchestration failures, especially in platforms such as Docker and Kubernetes (Watada et al., 2019). Chaos engineering tests on containers simulate abrupt stoppages and network disruptions, helping to evaluate containers’ ability to recover within container orchestration systems while maintaining application performance  (Simonsson et al., 2021; Camacho et al., 2022; Bremmers, 2021; Kostic, 2024; Ikeuchi et al., 2020).
Serverless functions allow applications to be triggered by specific events and run in isolated containers without requiring infrastructure management (Castro et al., 2017; Taibi et al., 2020). While serverless architectures are scalable and cost-effective, they face challenges such as resource limitations and the reliance on third-party cloud providers (Rajan, 2020; van Eyk and Iosup, 2018; Kadikar, 2023; Camacho et al., 2022). Chaos engineering helps ensure that serverless functions can handle these challenges by simulating cold starts and execution failures, verifying that functions recover quickly and maintain performance under varying loads (Starr et al., 2022; Jernberg et al., 2020; Gill, 2022; Bedoya et al., 2023).
Cloud
Cloud environments are integral to modern Information Technology (IT) infrastructures, providing flexible and scalable resources to meet dynamic business demands (Avram, 2014; Garrison and Nova, 2017). These environments are typically classified into three main models: public, private, and hybrid clouds (Balasubramanian and Aramudhan, 2012; Gorelik, 2013). Each cloud model has unique characteristics, risks, and opportunities, where chaos engineering can address potential failures.
Public cloud platforms, such as AWS, Microsoft Azure, and Google Cloud, offer computing resources over the Internet that are shared between multiple organizations (Saraswat and Tripathi, 2020). This model enables businesses to scale their infrastructure on demand, eliminating the need for significant investments in physical hardware. Due to its cost efficiency and scalability, the public cloud has become a popular choice for many enterprises (Dutta and Dutta, 2019). However, relying on shared infrastructure introduces specific vulnerabilities. For example, a regional outage or failure in an availability zone can have widespread impacts, simultaneously affecting many customers (Paavolainen, 2016). In addition, many organizations depend on third-party services, such as application programming interfaces (APIs) or managed databases, and failures in these services can trigger further disruptions in their systems (Zhou et al., 2010). To address these risks, chaos engineering can simulate large-scale failures, such as regional outages or availability zone failures, to test how well systems can failover to other regions without significant service interruptions (Fogli et al., 2023). Moreover, by injecting failures into critical third-party services, chaos engineering enables organizations to evaluate the effectiveness of their recovery strategies, ensuring that operations can continue even in the face of service disruptions (Ji et al., 2023; Yu et al., 2021).
Private cloud environments, which provide dedicated infrastructure exclusively for a single organization, can be hosted on-premises or managed by third-party providers, offering greater control over data privacy, security, and customization (Jadeja and Modi, 2012; Gastermann et al., 2015). This model suits organizations with strict regulatory requirements or managing sensitive information (Pearson, 2013). Despite their benefits, private clouds often face scalability and geographic redundancy challenges, as they typically lack the built-in redundancy of public clouds, making them more susceptible to service interruptions due to hardware failures or resource constraints (Kumar and Goudar, 2012). Without adequate failover mechanisms, private cloud environments may struggle to maintain operational continuity during hardware or network failures (Rimal et al., 2009; Tivey et al., 2019). Chaos engineering can be used in private clouds to simulate hardware failures, such as server crashes or network partitions, to evaluate the system’s ability to recover from localized disruptions (Patel, 2022). By simulating resource exhaustion, such as CPU or memory overloads, organizations can test whether their infrastructure can scale to meet unexpected surges in demand, ensuring continuous service availability under pressure (Gill, 2021).
Finally, hybrid cloud environments combine the strengths of public and private clouds, allowing organizations to manage sensitive workloads on private infrastructure while leveraging the scalability of public clouds for nonsensitive or high-demand tasks (Goyal, 2014). This approach provides flexibility but introduces complexity in maintaining smooth integration between the two environments (Lu et al., 2014). A key challenge of hybrid clouds is to ensure seamless communication between public and private components, as issues such as network latency, connectivity problems, and data synchronization delays can arise when workloads are distributed across both environments, potentially leading to service disruptions (Annapureddy, 2010). Chaos engineering can address this issue by allowing organizations to simulate network interruptions or increased latency between public and private clouds (Fogli et al., 2023). These tests help to evaluate whether workloads can continue to operate despite connectivity issues (Torkura et al., 2020). Furthermore, chaos experiments that simulate delays or failures in data replication help ensure that the system maintains data consistency and operational continuity, even when part of the infrastructure is compromised (Camacho et al., 2022).
On-Premise
This refers to the computing infrastructure hosted and managed within an organization’s physical location rather than relying on cloud providers (Willard and Hutson, 2024; Nikolaidis et al., 2021; Väisänen, 2023). These systems offer greater control over hardware, data, and network resources, essential for industries with strict regulatory requirements. However, they also face hardware failures and scaling limitations, affecting the reliability of the systems (Byström, 2022). For example, an organization managing critical on-premise data might simulate a hardware failure, such as a server or disk crash, to test how quickly backup systems respond. If synchronization delays are found, the replication process can be optimized to ensure smooth transitions and compliance with data availability standards (Fogli et al., 2023; Dedousis et al., 2023). Similarly, scaling limitations can be addressed by testing how an on-premise infrastructure responds to various demand surges, such as a sudden spike in user traffic during an e-commerce company’s high-demand event (Azure Readiness, 2023; Gianchandani et al., 2022). This helps identify bottlenecks, optimize resource allocation, and ensure that the system remains responsive under varying load conditions (Naqvi et al., 2022; Wu, 2021; Sbiai et al., 2023; Wickramasinghe, 2023).
5.1.2. Experimentation Modes
Chaos engineering platforms offer three modes of experimentation, namely manual, semi-automated, and fully automated, each tailored to different levels of control, complexity, and system scale.
Manual experimentation is most appropriate when engineers require high precision and control, as they take full responsibility for designing, configuring, and executing chaos tests (Kostic, 2024). This approach is suited for organizations with little or no experience in chaos engineering or when highly specific, customized experiments are necessary (Durai, 2022). For example, an engineer might manually trigger a database failure or introduce network latency to observe the system response in real time (Jernberg et al., 2020; Bremmers, 2021). While this mode allows for fine-tuned control, it can be time-consuming and prone to human error, making it ideal for small-scale, customized tests or critical systems where close monitoring is essential (Ikeuchi et al., 2020; Chen et al., 2024).
Semi-automated experimentation provides a middle ground between manual control and complete automation, where engineers define the experiment parameters (Sondhi et al., 2021; Jernberg et al., 2020). At the same time, the platform automates tasks such as triggering tests and collecting data, making this mode ideal for experiments that benefit from automated execution while still requiring moderate human involvement to reduce manual effort (Kesim et al., 2020; Nikolaidis et al., 2021; Gill, 2021). For example, engineers might set up an experiment to test resource exhaustion on specific microservices, with the platform automatically running the test at predefined intervals, collecting performance metrics, and allowing them to focus on adjusting parameters and interpreting results (Pierce et al., 2021; Dedousis et al., 2023).
Fully automated experimentation suits large-scale systems that require frequent and continuous testing (Zhang et al., 2021a; Simonsson et al., 2021; Basiri et al., 2019). After engineers configure the initial setup, the chaos engineering platform takes over, seamlessly integrating experiments into the system operations, typically through the CI/CD pipelines (Konstantinou et al., 2021; Pulcinelli et al., 2023; Manna, 2021; Colyer, 2019; Azure Readiness, 2023). For example, after a code deployment or infrastructure update, the platform can automatically simulate a service outage or introduce network disruptions, then autonomously gather performance metrics, identify system weaknesses, and generate detailed reports, requiring engineers to intervene only when necessary (Dedousis et al., 2023; Bairyev, 2023; Kesim et al., 2020; Cotroneo et al., 2022; Nikolaidis et al., 2021; Patel, 2022).
5.1.3. Chaos Automation Strategy
Chaos engineering platform employs an automation strategy to ensure consistent and repeatable experiments, reduce manual effort, and enable continuous testing (Kostic, 2024; Wickramasinghe, 2023; Inc., 2023).
Custom Code
This involves developing programs or scripts to extend the capabilities of existing tools and gain precise control over system configurations and experiments (Myers et al., 2000; Thüm et al., 2014; Khankhoje, 2022). It allows engineers to simulate unique failure conditions, ranging from simple component failures, such as restarting a single service, to complex multi-step sequences involving multiple systems, tailored to the specific challenges of their architecture that standard tools may not address (Gill, 2021; Inc., 2023; Torkura et al., 2020; Azure Readiness, 2023; Patel, 2022; Nikolaidis et al., 2023). For example, custom code can terminate random containers in distributed environments such as Kubernetes, simulating unexpected service failures and testing the system’s recovery capacity by automatically restarting services (Bremmers, 2021; Fawcett, 2020; Simonsson et al., 2021; Vu et al., 2022). In more complex scenarios, custom code can initiate a service outage followed by a database failure or a network partition to observe how these failures impact system recovery and resilience (Wickramasinghe, 2023; Gill, 2021; Simonsson et al., 2021). Additionally, custom code can be integrated into CI/CD pipelines to simulate high resource consumption, such as CPU throttling or memory exhaustion, after each deployment, allowing early detection of performance bottlenecks before they affect production (Azure Readiness, 2023; Konstantinou et al., 2021; Manna, 2021; Kadikar, 2023; Palani and Gupta, 2023). This custom code for chaos experiments can be written in various programming languages, such as Python, Bash, or Go, depending on the complexity and specific requirements of the experiment (Siwach et al., 2022; Malik et al., 2023; Sondhi et al., 2021; Simonsson et al., 2021; Zhang et al., 2021b).
Auto-Integration of Third-Party Tools
This process automatically integrates external systems, such as monitoring, log, observability, alerting, incident management, and security platforms, into chaos engineering experiments (Torkura et al., 2020; Simonsson et al., 2021). This automated integration ensures continuous data collection without manual intervention, enabling real-time analysis during system disruptions or failures. As a result, teams can monitor performance, identify issues quickly, and respond more efficiently to failures (Palani and Gupta, 2023; Kostic, 2024). This process is vital to integrating monitoring tools that track essential system metrics such as CPU usage, memory consumption, and network performance (Abdul, 2024; Wu, 2021; Gianchandani et al., 2022; Zhang et al., 2019). Monitoring tools can automatically detect performance deviations and provide immediate feedback. For example, if a service is overloaded due to an unexpected spike in user requests, monitoring systems measure the increase in CPU usage and network load, helping the team assess the impact and determine whether the system can handle the demand or if further intervention is needed (Wickramasinghe, 2023; Fogli et al., 2023; Cotroneo et al., 2022; Gill, 2021; Siwach et al., 2022; Malik et al., 2023). Logging tools record vital events such as service failures, error messages, and recovery attempts, documenting the system’s response in real-time. For example, if network communication between two microservices fails, the logging systems capture the exact moment the failure occurred and show the error messages generated by failed retries (Cotroneo et al., 2022; Ma et al., 2023). This data helps teams analyze how the failure affected communication and whether recovery mechanisms worked effectively (Naqvi et al., 2022; Malik et al., 2023; Kostic, 2024; Palani and Gupta, 2023). Building on this, observability tools provide a broader view by correlating logs and metrics to show how different components of the system interact during failure scenarios. These tools enable teams to identify how failures propagate across interconnected services (Simonsson et al., 2021; Manna, 2021; Green, 2023; Kostic, 2024; Krishnamurthy, 2021). For example, if latency is introduced between services in a chaos experiment, observability systems track the delays at each hop between services and how this latency affects downstream components such as database queries or API responses. This tracing allows the team to see if minor delays in one service cause issues elsewhere (Durai, 2022; Katirtzis, 2022; Kostic, 2024; Chen et al., 2024; Nikolaidis et al., 2023; Frank et al., 2023). Furthermore, alert notification systems can ensure that relevant teams are promptly notified when critical failures occur or performance thresholds are exceeded (Tivey et al., 2019; Gianchandani et al., 2022; Palani and Gupta, 2023; Bremmers, 2021; Starr et al., 2022). For example, if a database connection pool reaches its maximum capacity during a high-load chaos test, the alert system triggers a notification to the database operations team, allowing them to address the issue before it causes a system-wide failure. This early warning helps prevent more significant disturbances (Green, 2023; Malik et al., 2023). Finally, incident management tools streamline the resolution process in conjunction with alerting by automatically generating tickets for failures detected during chaos tests. These tickets are assigned to the appropriate teams and tracked until the issue is resolved (Inc., 2023; Patel, 2022; Green, 2023; Le, 2022). For example, if a chaos experiment causes a loss of connection to external APIs, the incident management system creates a ticket assigned to the appropriate API support team, ensuring that the problem is tracked and resolved efficiently (Rivera et al., 2023; Siwach et al., 2022).
Workflow
A chaos engineering workflow can automate the controlled execution of experiment steps, such as setting up the environment, executing failures, monitoring, and cleanup, ensuring that the experiments are consistent and repeatable  (Camilli et al., 2022). This reduces the risk of errors and guarantees the reliability in each run. For example, automating a database outage simulation ensures that critical steps such as downgrading the database, monitoring the system response, and restoring functionality are executed reliably every time  (Wickramasinghe, 2023). In addition, workflows simplify complex experiments that involve multiple systems or services. In such cases, they can coordinate the simultaneous failure of various microservices, allowing teams to test the system’s ability to maintain service availability despite several disruptions  (Nikolaidis et al., 2021; Manna, 2021). Furthermore, a workflow engine ensures that each step is executed correctly while capturing and analyzing the results, making the experimentation process both structured and efficient (Torkura et al., 2020; Bairyev, 2023).
Prioritization
This involves categorizing system components into critical and non-critical groups to focus testing efforts on the areas where failures would have the greatest impact (Dedousis et al., 2023; Anwar and Arunachalam, 2019). Critical components are essential for maintaining core operations, and their failure would cause significant disruption (Fogli et al., 2023; Kesim et al., 2020). For example, API gateways and distributed databases are critical because these failures can disrupt communication between services, cause widespread outages, or lead to data loss and downtime, directly affecting users (Siwach et al., 2022). In contrast, non-critical components have minimal immediate impact on the core functionality of the system. For instance, log aggregation systems are non-critical; their failure may affect internal monitoring but will not disrupt user-facing services or essential system functions (Dedousis et al., 2023). By prioritizing critical components, testing can focus on areas where failure would result in the most severe consequences. In contrast, the resilience issues of non-critical components can be addressed less urgently, since their failure poses a lower risk to overall system stability (Wickramasinghe, 2023).
5.1.4. Deployment
The type of deployment selected for chaos experiments is a critical decision in chaos engineering, as it defines the testing environment and significantly affects the performance and stability of the system (Meiklejohn et al., 2021; Abdul, 2024; Zhang et al., 2019). Choosing the appropriate deployment type ensures that the experiments yield meaningful insights while minimizing the risks to live operations (Siwach et al., 2022; Wickramasinghe, 2023). This section identifies two primary deployment types: production and pre-production.
Production.
This refers to running tests or experiments in a live environment where real users actively interact with the System Under Test (SUT) (Zhang et al., 2019; Bremmers, 2021; Vu et al., 2022; Nikolaidis et al., 2023; Cotroneo et al., 2022). For example, deploying a new feature on an e-commerce website during sales periods allows teams to observe user interactions and system performance under actual conditions (Jernberg et al., 2020). A significant benefit of production deployment is the ability to uncover vulnerabilities that might remain hidden in more isolated testing environments (Wickramasinghe, 2023; Nikolaidis et al., 2021). For example, testing a service during peak traffic can reveal weaknesses that would not otherwise emerge, making it crucial to gain insight into how the SUT handles real-world disruptions and help identify areas for further optimization (Fogli et al., 2023; Chen et al., 2024; Simonsson et al., 2021). Production deployment introduces risks, as failures during testing can directly affect users and disrupt business operations (Roa, 2022; Jernberg et al., 2020). Production experiments must be carefully planned to mitigate these risks, such as scheduling tests during low-traffic periods and implementing automated rollback mechanisms to restore normal operations if problems occur (Sondhi et al., 2021; Bremmers, 2021).
Pre-production.
This environment closely mirrors the production environment, allowing teams to validate system functionality and performance before going live (Naqvi et al., 2022; Vu et al., 2022; Bremmers, 2021). An example of a pre-production environment is a staging environment, where the SUT can be deployed with the same configurations as production, enabling comprehensive checks on how it will perform under expected loads (Jernberg et al., 2020; Manna, 2021). Pre-production deployment allows for complex tests that might be too risky to execute in live environments (Bhaskar, 2022). For example, simulating a network partition can reveal how well the SUT responds to such failures, helping teams resolve potential problems before they reach production (Green, 2023). This added layer of testing ensures major releases and helps reduce the risk of failure once the system is live  (Kostic, 2024). One limitation of pre-production testing is that it may not accurately reflect the complexities of real-world operations (Bhaskar, 2022; Naqvi et al., 2022). Consequently, these tests should be viewed as a complement, rather than a substitute, to production testing (Zhang et al., 2019; Torkura et al., 2020; Gianchandani et al., 2022).
5.1.5. Evaluation Approaches
Evaluation refers to determining the effectiveness of chaos engineering experiments by measuring their impact on system stability, the system’s ability to handle induced failures, and the growth and maturity of the organization’s chaos engineering practices. Evaluation approaches can be categorized into qualitative and quantitative methods.
Quantitative Evaluation
This involves systematically measuring numerical data to assess a system’s performance under failure conditions. It focuses on metrics that can be measured, such as response time, error rates, and system availability (Garg et al., 2013). By comparing the behavior of the system during chaos experiments with its regular operation, engineers can identify weaknesses and areas for improvement based on objective data. We identified eight quantitative evaluation approaches, which include the following:
  • Response Time Metrics measure the total time it takes for a system to process a request, from when it is received to when the result is returned, encompassing both processing and response delivery (Frank et al., 2021a; Abdul, 2024; Frank et al., 2021b; Dua, 2024; Cotroneo et al., 2022; Green, 2023; Kostic, 2024; Mooney, 2023; Siwach et al., 2022; Camacho et al., 2022; Sondhi et al., 2021; National Australia Bank, 2020; Al-Said Ahmad et al., 2024; Kesim et al., 2020). For example, a baseline response time of 18 milliseconds might increase to 27 milliseconds during chaos experiments such as ’KillNodes’, where parts of the system are intentionally shut down to test disruptions, indicating slower request processing and completion (Torkura et al., 2021).
  • Latency Metrics measure the initial delay or waiting time between when a request is made and when the system begins processing that request (Fogli et al., 2023; Wickramasinghe, 2023; Kostic, 2024; Poltronieri et al., 2022; Bremmers, 2021; Cotroneo et al., 2022; Azure Readiness, 2023; Colyer, 2019). It does not include the total processing time, only the delay before the system starts responding (Cotroneo et al., 2022; Durai, 2022; Fawcett, 2020; Gianchandani et al., 2022; Frank et al., 2021a; Siwach et al., 2022; Yu et al., 2021). For example, latency can increase from 44 to 70 milliseconds during chaos experiments, signaling that the system takes longer to react to requests (Kassab et al., 2022).
  • Resource Utilization Metrics track the consumption of system resources, such as CPU and memory (Azure Readiness, 2023; Fogli et al., 2023; Torkura et al., 2020; Simonsson et al., 2021; Zhang et al., 2021a; Camacho et al., 2022; Pulcinelli et al., 2023; Wu, 2021; Gianchandani et al., 2022; Roa, 2022; Kadikar, 2023; Green, 2023; Zhang et al., 2019). For example, CPU usage may spike from 50% to 80% during chaos experiments, indicating an increased strain on the system due to failures  (Simonsson et al., 2021).
  • Error Rate Metrics measure the frequency of errors encountered during system operation (Gill, 2021; Zhang et al., 2021a; Cotroneo et al., 2022; Basiri et al., 2019; Rivera et al., 2023; Roa, 2022; Kostic, 2024; Vinisky, 2024; Colyer, 2019; Durai, 2022). For instance, an increase in the error rate from 0.5% to 5% during chaos experiments suggests that the system is struggling to handle failures effectively (Zhang et al., 2023).
  • Availability Metrics refers to a system’s ability to remain operational, measured by uptime and downtime (Torkura et al., 2020; Gill, 2021; Camacho et al., 2022; Nikolaidis et al., 2021; Ma et al., 2023). Uptime indicates how long the system is fully functional, while downtime shows when the system is unavailable (Shah, 2021; Durai, 2022; Abdul, 2024; Gianchandani et al., 2022).
  • Throughput Metrics measure the volume of data or transactions the system can handle in a given time (Kostic, 2024; Vinisky, 2024; Al-Said Ahmad et al., 2024; Torkura et al., 2020). For example, throughput might drop from 1,000 transactions per second to 600 during chaos experiments, leading to delays in processing user requests or data transfers (Kassab et al., 2022; Abdul, 2024).
  • Mean Time to Recovery (MTTR) measures the average time required to restore the system after a failure (Durai, 2022; Abdul, 2024; Starr et al., 2022; Gianchandani et al., 2022). A higher MTTR, indicating slower recovery, reveals areas that need faster recovery processes (Starr et al., 2022; Frank et al., 2021b; Camacho et al., 2022; Bhaskar, 2022). For example, if a cloud-based e-commerce platform experiences a server crash during a flash sale and MTTR increases from 2 to 10 minutes, it signals the need for more efficient recovery mechanisms to prevent revenue loss and customer frustration (Siwach et al., 2022).
  • Mean Time Between Failures (MTBF) refers to the average time that a system operates without failure, reflecting the reliability of the system (Bhaskar, 2022; Camacho et al., 2022). A lower MTBF suggests frequent failures, highlighting the need for improved stability (Kadikar, 2023). For example, if a healthcare system that manages patient records experiences failures every two hours, it points to the need for a more reliable system architecture to ensure uninterrupted access to patient data during critical procedures.
Qualitative Evaluation
This gathers non-numerical insights to assess system performance during failures, focusing on user feedback, team experiences, and business impacts. By understanding these aspects, engineers gain deeper insight into the broader effects of disruptions. This complements quantitative methods by adding meaning to numerical data. We identified four qualitative evaluation approaches, which include:
  • User Experience Evaluation involves gathering qualitative feedback on how system incidents impact end users (Gill, 2021; Ji et al., 2023). This may include conducting surveys, interviews, or usability tests to assess how failures affect users’ ability to interact with the system (Camacho et al., 2022; Kostic, 2024). For example, users might report frustration or difficulty in completing tasks during a chaos experiment that causes slow page load times or feature outages (Fawcett, 2020; Sbiai et al., 2023; Gill, 2022). This evaluation is crucial because it reveals how technical failures affect user satisfaction and retention, providing insight beyond response time or availability figures (Wu, 2021; Le, 2022; Cotroneo et al., 2022).
  • Business Impact Evaluation assesses how chaos experiments affect key business metrics, including revenue, customer churn, and operational efficiency (Inc., 2023; Jernberg et al., 2020). This involves understanding specific financial losses, customer complaints, and service-level breaches resulting from experiments(National Australia Bank, 2020; Palani and Gupta, 2023). For instance, a chaos experiment that causes downtime during a product launch could result in significant lost sales and a spike in customer service tickets (Ji et al., 2023; Starr et al., 2022). This evaluation is critical to linking technical failures with direct business consequences, enabling more targeted investment in stability measures where it matters most (Manna, 2021; Moon, 2022).
  • System Behavior and Observability focuses on how effectively system components behave and are monitored during chaos experiments (Anwar and Arunachalam, 2019; Yu et al., 2021; Meiklejohn et al., 2021; Mooney, 2023). It involves reviewing logs/traces and monitoring gaps to identify areas where failures are not easily detected or diagnosed (Ikeuchi et al., 2020; Soldani and Brogi, 2021; Zhang et al., 2019; Green, 2023). For example, an experiment might expose missing alerts or inaccurate telemetry data for critical services, making it harder for engineers to respond (Colyer, 2019; Meiklejohn et al., 2021; Basiri et al., 2019). This evaluation helps to ensure that the system’s observability tools are robust enough to detect problems early and aid in faster diagnosis and recovery (Dedousis et al., 2023; Fawcett, 2020).
  • Team Feedback captures insights from engineering teams on their ability to respond to failures during chaos experiments (Kostic, 2024; Fogli et al., 2023; Torkura et al., 2020). It evaluates the effectiveness of incident response, communication, and decision-making processes under system stress (Zhang et al., 2019; Simonsson et al., 2021). For example, feedback might highlight confusion in handling specific failure scenarios or delays in coordinating recovery efforts (Inc., 2023; Krishnamurthy, 2021; Chen et al., 2022). This evaluation is essential to identify operational weaknesses in real-time incident management and to improve team workflows and readiness for actual outages (Kesim et al., 2020; Wu, 2021; Fawcett, 2020).
5.2. Chaos Engineering Tools
We used the proposed taxonomy to evaluate 10 popular chaos engineering tools (T1-T10 in Table 10). We selected these tools by identifying the most-cited tools from our pool of reviewed papers, then ranking them using GitHub metrics, precisely the ’Most stars’ measure, to confirm their popularity and community engagement (Borges et al., 2016). Collectively, these tools had over 93,000 stars as of November 2024. These tools have been mapped according to their applicability to different execution environments, automation modes, strategies, and deployment stages. This categorization enables practitioners to make well-informed choices by aligning tool selection with specific infrastructure needs, operational requirements, automation capabilities, and system risk tolerance. Consequently, selecting the appropriate tool in a software development project becomes a crucial decision (Kavaler et al., 2019). For example, T1 and T2 provide manual and semi-automated chaos testing capabilities, supporting Kubernetes, Docker, and virtual machine setups. While T1 supports performing chaos experiments in live production environments, T2 focuses on testing environments, particularly pre-production setups. T3 is a retired tool designed initially for AWS but lacks support for Kubernetes. Its functions have been migrated to newer Netflix tools such as Chaos Monkey, Swabbie, and Spinnaker, making it better suited to traditional infrastructures. T4 and T5 both support Kubernetes and virtual machine environments in hybrid clouds. These tools offer full automation options, with features like workflow automation and custom integrations, making them ideal for chaos testing in pre-production and production environments. T6 and T7 focus on Kubernetes and serverless functions, offering manual, semi-automated, and fully automated modes. While T6 is optimized for hybrid cloud setups with advanced workflow customization, T7 is better suited for production resilience testing in Kubernetes environments. T8 and T9 emphasize adaptability across diverse infrastructures. T8 supports Kubernetes, serverless functions, and on-premise setups, focusing on manual testing modes with customizable workflows. T9, on the other hand, operates across virtual machines and Kubernetes environments, providing semi-automated and fully automated capabilities with workflow automation and custom integration options, making it suitable for public cloud systems. Finally, T10 excels in hybrid cloud environments, supporting Kubernetes and virtual machines with advanced automation and integration capabilities for pre-production and production stages. This taxonomy ensures that chaos engineering tools are mapped to their ideal use cases, helping practitioners choose the best tools for their systems. Our online appendix provides the complete list of tools considered.
Figure 9. A Taxonomy of Chaos Engineering Platforms.
[…]
A taxonomy of chaos engineering platforms.
Table 10. Top-10 Most Used Chaos Engineering Platforms.
Tool Code Tool Name Execution Environment Automation Mode Automation Strategy Deployment Stages
Endpoint On- Premise Cloud Manual Semi- Automated Fully Automated Prioritization Workflow Automation Custom Code Auto- Integration Pre- production Production
Docker Kubernetes
Virtual
Machine
Serverless
Function
Private Public Hybrid
T1 Chaos Monkey
T2 Toxiproxy
T3 Simian Army
T4 Chaos Mesh
T5 Chaosblade
T6 Litmus
T7 Pumba
T8 Chaos Toolkit
T9 Powerful seal
T10 Chaoskube
5.3. Adoption Practices
Adopting chaos engineering requires a structured approach, balancing best practices with avoiding common mistakes. To effectively implement chaos experiments and reduce the risk of system failures or disruptions, organizations should follow well-established guidelines for planning, executing, and evaluating these experiments  (Siwach et al., 2022; Green, 2023). Table 11 summarizes essential best practices, such as starting with small, controlled experiments, setting clear objectives, automating tests, and closely monitoring system behavior. Additionally, organizations should limit the scope of experiments, encourage cross-team collaboration, and continuously refine processes based on experiment outcomes. While following best practices is crucial, organizations must also avoid common pitfalls, such as those outlined in Table 12, which include failing to tailor experiments to specific systems, neglecting proper planning and hypothesis formation, and lacking adequate safety mechanisms. By adhering to best practices and avoiding these common pitfalls, organizations can confidently integrate chaos engineering into their systems and achieve successful outcomes (Siwach et al., 2022).
Table 11. Best Practices
No Best Practices Description Articles Total
1 Start Small and Scale Up Start with manageable tests, then expand gradually. (Fogli et al., 2023; Torkura et al., 2020; Bremmers, 2021; Patel, 2022; Gill, 2021; Chen et al., 2022; Siwach et al., 2022; Jernberg et al., 2020; Anwar and Arunachalam, 2019; Palani and Gupta, 2023; Kadikar, 2023; Green, 2023) 12
2 Define Clear Objectives Set precise goals to guide each experiment’s purpose and outcomes. (Torkura et al., 2020; Dedousis et al., 2023; Inc., 2023; Patel, 2022; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Kesim et al., 2020; Soldani and Brogi, 2021; Siwach et al., 2022; Jernberg et al., 2020; Jakkaraju, 2020; Dua, 2024; Kadikar, 2023; Green, 2023) 14
3 Automate Experiments Use automation tools to run chaos experiments regularly, ensuring consistency and frequency. (Torkura et al., 2020; Zhang et al., 2019; Inc., 2023; Gill, 2021; Wickramasinghe, 2023; Camacho et al., 2022; Zhang et al., 2021a; Al-Said Ahmad et al., 2024; Vu et al., 2022; Kassab et al., 2022; Ikeuchi et al., 2020; Naqvi et al., 2022; Malik et al., 2023; Soldani and Brogi, 2021; Siwach et al., 2022; Jernberg et al., 2020; Bhaskar, 2022; Gill, 2022; Fawcett, 2020; Bairyev, 2023; Sbiai et al., 2023; Azure Readiness, 2023; Roa, 2022; Green, 2023) 24
4 Track and Measure Results Monitor the system’s performance during experiments and collect data to see the impact. (Fogli et al., 2023; Torkura et al., 2020; Dedousis et al., 2023; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Camacho et al., 2022; Torkura et al., 2021; Zhang et al., 2021a; Vu et al., 2022; Kassab et al., 2022; Ikeuchi et al., 2020; Malik et al., 2023; Soldani and Brogi, 2021; Siwach et al., 2022; Konstantinou et al., 2021; Jernberg et al., 2020; Hawkins, 2020; Manna, 2021; Azure Readiness, 2023; Kadikar, 2023; Green, 2023) 23
5 Involve all Key Stakeholders Include team members to ensure everyone is informed and aligned. (Torkura et al., 2020; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Frank et al., 2021b; Siwach et al., 2022; Dua, 2024; Kadikar, 2023; Green, 2023; Gill, 2022; Kostic, 2024) 12
6 Document and Improve Record experiments to continually improve the system. (Torkura et al., 2020; Zhang et al., 2019; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Siwach et al., 2022; Palani and Gupta, 2023; Kadikar, 2023; Green, 2023; Kesim et al., 2020) 11
Table 12. Bad Practices
No Bad Practices Description Articles Total
1 Misaligned Chaos Chaos experiments that don’t match the system yield misleading results. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Vu et al., 2022; Nikolaidis et al., 2023; Ma et al., 2023; Sondhi et al., 2021; Jernberg et al., 2020; Konstantinou et al., 2021; Park et al., 2019; Zhang et al., 2023; Rivera et al., 2023; Patel, 2022; Le, 2022; Lardo, 2019) 16
2 Inadequate Experimentation Without planning, clear goals, and baseline tracking, assessing chaos test impacts becomes difficult. (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Jakkaraju, 2020; Le, 2022; Hawkins, 2020; Anwar and Arunachalam, 2019; Dua, 2024; National Australia Bank, 2020; Mooney, 2023; Starr et al., 2022; Palani and Gupta, 2023; Gill, 2022; Fawcett, 2020; Gianchandani et al., 2022; Bairyev, 2023; Abdul, 2024; Azure Readiness, 2023; Kadikar, 2023; Green, 2023; Aharon, 2024; Krivas and Portela, 2020; Kostic, 2024; Frank et al., 2021b; Jernberg et al., 2020; Konstantinou et al., 2021; Ji et al., 2023; Soldani and Brogi, 2021; Park et al., 2019; Wachtel, 2023) 31
3 Safety Mechanism Oversight Not testing rollback and safety features risks uncontrolled failures. (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Patel, 2022; Jakkaraju, 2020; Bhaskar, 2022; Katirtzis, 2022; Dua, 2024; Starr et al., 2022; Tivey et al., 2019; Palani and Gupta, 2023; Bairyev, 2023; Sbiai et al., 2023; Kadikar, 2023; Green, 2023; Vinisky, 2024; Pierce et al., 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Ikeuchi et al., 2020; Chen et al., 2022; Naqvi et al., 2022; Soldani and Brogi, 2021; Konstantinou et al., 2021) 25
4 Insufficient Tooling Lacking observability, automation, and integration tools makes monitoring and analyzing chaos experiments difficult. (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Hawkins, 2020; Starr et al., 2022; National Australia Bank, 2020; Tomka, 2024; Dua, 2024; Mooney, 2023; Lardo, 2019; Gianchandani et al., 2022; Abdul, 2024; Azure Readiness, 2023; Roa, 2022; Kadikar, 2023; Green, 2023; Krivas and Portela, 2020; Kostic, 2024; Simonsson et al., 2021; Nikolaidis et al., 2021; Soldani and Brogi, 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Naqvi et al., 2022; Malik et al., 2023; Camilli et al., 2022; Bremmers, 2021; Patel, 2022; Le, 2022; Zhang et al., 2021a; Cotroneo et al., 2022; Ikeuchi et al., 2020; Bedoya et al., 2023; Basiri et al., 2019; Pulcinelli et al., 2023; Jernberg et al., 2020; Anwar and Arunachalam, 2019; Durai, 2022; Jakkaraju, 2020) 39
5 Poor Stakeholder Collaboration Excluding key team members causes misalignment and resistance. (Bremmers, 2021; Starr et al., 2022; Le, 2022; Krishnamurthy, 2021; Katirtzis, 2022; Durai, 2022; Tivey et al., 2019; Gill, 2022; Green, 2023; Poltronieri et al., 2022; Frank et al., 2021b; Al-Said Ahmad et al., 2024; Jernberg et al., 2020; Kesim et al., 2020) 14
6. Current State of Chaos Engineering (RQ4)
This section first presents the classification of research in chaos engineering, then examines the number of sources published per year and venue type, and finally reviews the venue rankings and key contributors in the field.
6.1. Number of Sources by Research Type
To better understand the current state of chaos engineering research, we classified the primary studies based on the research classification framework proposed by Wieringa et al. (Wieringa et al., 2006). This framework includes the following categories:
  • Evaluation research: These studies investigate problems in software engineering practice, often uncovering new knowledge about causal or logical relationships among different phenomena. This type of research is conducted through case or field studies, field experiments, and surveys.
  • Philosophical papers: These studies introduce new conceptual frameworks or ways of thinking.
  • Solution proposal papers: These studies suggest techniques and argue for their relevance without full-scale validation.
  • Validation research: These studies rigorously examine the properties of a proposed solution that has yet to be implemented in practice, employing methodologically sound approaches such as experiments and simulations.
  • Opinion papers: These studies present the author’s views on what is good or bad about a particular issue and how things should be approached.
  • Personal experience papers: These studies describe what was done based on the author’s own experience with one or more projects rather than explaining why it was done.
The first author did the classification, which was then reviewed by the second author, with all discrepancies resolved through discussion. The results of this classification, shown in Figure 12, indicate that solution proposal papers were the most frequent type of research between 2019 and 2024. In the academic literature, solution proposal papers dominated, with nine papers in 2021 and 2022 and slightly fewer (eight) in 2023. The grey literature, on the other hand, showed a marked rise in experience reports, growing from six reports in 2022 to eight in 2023. This classification highlights research activities’ evolving distribution and focus, emphasizing the dominance of solution proposals in academic literature and the increasing role of practical experience reports in grey literature.
6.2. Number of Sources by Year and Venue Type
We plot the distribution of chaos engineering studies from 2019 to 2024 to examine its adoption pattern (see Figure 10). In 2019, academic publications were minimal, limited to conferences and workshops, while grey literature focused on blogs, reflecting an early exploratory phase. By 2020, academic interest grew, focusing on conferences, though grey literature centered around blogs. In 2021, academic publications peaked across journals, conferences, and workshops, marking the “early adopter” phase, while grey literature remained stable, illustrating the challenges of Moore’s Chasm (see Figure 1). In 2022, academic output slightly declined. Still, grey literature grew, diversifying with blogs and white papers, suggesting chaos engineering was beginning to bridge Moore’s Chasm and gain broader acceptance. By 2023, academic publications remained steady, though workshops declined, while grey literature expanded further, especially with white papers, signaling continued industry engagement. The slight decline observed in 2024 for both academic and grey literature suggests a possible temporary slowdown, as the year is not yet complete, and more publications may emerge (Werbińska-Wojciechowska and Winiarska, 2023).
A precise distribution emerges between academic and grey literature publications when examining venue types. Of the 44 scholarly publications, 28 (63.64%) were conference papers, 11 (25%) were journal articles, and 5 (11.36%) were workshops, reflecting a strong preference for conferences as the primary channel for academic dissemination. Among the 44 grey literature publications, 41 (93.18%) were blogs, while 3 (6.82%) were white papers, underscoring blogs as the dominant platform for informal knowledge sharing. These trends indicate that chaos engineering, which began as an emerging research area, has evolved into a mature field of interest (Varela-Vaca and Quintero, 2021), attracting significant attention from academia and industry.
Figure 10. Distribution of Academic and Grey Literature Studies over Years and Venue Type.
[…]
Distribution of academic(N = 44) and grey(N = 44) literature studies over years and venue type.
6.3. Publication trends and key Contributors in Academic and Grey Literature
We analyzed publication trends in chaos engineering by examining venues based on their type and ranking and identifying key contributors within academic and grey literature. This analysis provides insights into the maturity of the domain (Varela-Vaca and Quintero, 2021).
6.3.1. Academic Literature
First, we analyzed venues based on their name, ranking, and type (journal or conference). As shown in Table 13, prominent venues include key publishers such as IEEE and ACM, which are widely recognized for hosting high-impact conferences and journals in chaos engineering. The venues were categorized using the Computing Research and Education Association of Australasia (CORE) ranking for conferences (CORE 2023) [^3] https://portal.core.edu.au/conf-ranks/ and the Scimago ranking for journals (Scimago 2023) [^4] https://www.scimagojr.com/journalrank.php. Figure 12 presents the distribution of sources in these venues. The data indicates that the largest share of studies is found in unranked venues (12 sources), with a significant number published in Q1 venues reflecting solid academic interest. High-ranking venues, such as A and A*, include several prominent IEEE and ACM conferences, emphasizing the quality of research in the field. Lower-ranked venues, including B, C, and workshops, also contribute to the research landscape, although fewer studies are published in Q2 venues. Table 13 highlights the top 10 primary studies in chaos engineering, ranked by citation count. It should be noted that only the first author of each study is listed, even if there are multiple authors  (Varela-Vaca and Quintero, 2021). Study A42 from Sun Yat-Sen University, China, leads the list, presenting MicroRank, a tool for end-to-end latency issue localization in microservice environments. Following closely are studies A02 and A09 from the Hasso Plattner Institute in Germany, which explore chaos engineering applications in cloud security and continuous auditing and threat detection across multi-cloud environments. KTH Royal Institute of Technology, Sweden, is represented by three influential works, A03, A05, and A06, investigating fault injection and resilience in containerized systems and Java Virtual Machine (JVM) environments. Netflix contributes with study A36, focusing on automating chaos experiments within production environments, while Carnegie Mellon University’s A22 emphasizes service-level fault injection testing via the Filibuster tool. Additional contributions include A26 from Lund University, which presents an implementation framework for chaos engineering, and A35 from the University of Stuttgart, which proposes an approach for prioritizing chaos experiments using established risk analysis techniques. This collection of studies from institutions in countries such as China, Germany, Sweden, and the United States reflects a global commitment to advancing chaos engineering with applications to enhance system resiliency, strengthen cloud infrastructure, and improve reliability in production environments worldwide.
6.3.2. Grey Literature
Unlike traditional academic literature, which follows a structured format and is ranked using systems such as h-index or Scimago, grey literature lacks standardized organization. It cannot be assessed using typical academic metrics. Instead, alt-metrics (alternative metrics used to measure online attention and engagement), such as comments, likes, shares, and social engagement, are used to gauge the quality and impact of grey literature. Table 14 highlights the top five articles in grey literature based on these alt-metrics. Grey literature on chaos engineering is often published by well-known IT companies (e.g., Microsoft, Gremlin, and Capital One) on company blogs or online publishing platforms (e.g., LinkedIn and Medium). These platforms are essential venues for sharing industry-driven insights in formats such as technical articles, blog posts, and white papers. The alt-metrics used, such as likes, comments, and shares, help measure the influence and engagement of the content within the industry, indicating its impact on chaos engineering. In cases where social engagement data was limited, additional quality assessments were considered, including the reputation of the venues, the authors’ expertise, and the clarity of the content (see Table  3).
As shown in Table 14, the top five grey articles are authored by industry professionals who provide valuable insights into chaos engineering. Key contributors come from prominent organizations, including Lyft, National Australia Bank (NAB), Expedia, Chaos Toolkit, and Eat Takeaway.com. For example, the study G12 from Lyft, published on Medium, received 348 likes and 2 comments for its introduction of a chaos experimentation framework. The study G11 from NAB, also published on Medium, received 156 likes for its focus on enhancing observability in microservices during chaos experiments. Additional notable contributions include the study G15 from Expedia, whose article on Medium earned 128 likes for detailing Kubernetes-integrated chaos engineering frameworks. The study G25, authored on LinkedIn by Chaos Toolkit, received 52 likes and 52 comments, emphasizing the importance of human interaction in chaos experiments. Lastly, the study G14 from Just Eat Takeaway.com, published on Medium, received 24 likes and 1 comment, highlighting the discovery of hidden system vulnerabilities through chaos engineering. These industry contributions have significantly influenced the development of chaos engineering practices, providing practical insights directly from leaders in the field. As discussed earlier, the same quality assessment criteria were applied, focusing on venue reputation, author expertise, and content clarity, when data on social engagement were insufficient (see Table  3).
Table 13. Top-10 of Papers Ordered by Citations in Academic Literature (SC - SCOPUS, GS - Google Scholar).
Rank ID Title Author Affiliation SC GS
1 A42 MicroRank: End-to-End Latency Issue Localization with Extended Spectrum Analysis in Microservice Environments Guangba Yu Sun Yat-Sen University, China 50 72
2 A02 CloudStrike: Chaos Engineering for Security and Resiliency in Cloud Infrastructure Kennedy A. Torkura Hasso Plattner Institute, Germany 39 70
3 A09 Continuous Auditing and Threat Detection in Multi-cloud Infrastructure Kennedy A. Torkura Hasso Plattner Institute, Germany 28 53
4 A03 A Chaos Engineering System for Live Analysis and Falsification of Exception-Handling in the JVM Long Zhang KTH Royal Institute of Technology, Sweden 15 44
5 A05 Observability and Chaos Engineering on System Calls for Containerized Applications in Docker Long Zhang KTH Royal Institute of Technology, Sweden 17 40
6 A36 Automating Chaos Experiments in Production Ali Basiri Netflix in Los Gatos, California. 15 40
7 A26 Getting Started with Chaos Engineering - Design of an Implementation Framework in Practice Hugo Jernberg Lund University and ICA Gruppen AB, Lund/Solna, Sweden 7 21
8 A22 Service-level Fault Injection Testing Christopher S. Meiklejohn Carnegie Mellon University, Pittsburgh, PA, United States 14 19
9 A06 Maximizing Error Injection Realism for Chaos Engineering With System Calls Long Zhang KTH Royal Institute of Technology, Sweden 6 18
10 A35 Identifying and Prioritizing Chaos Experiments by Using Established Risk Analysis Techniques Dominik Kesim University of Stuttgart, Germany 8 16
Table 14. Top-5 Articles Ordered by Altmetrics (Comments / Likes/Claps) in Grey Literature.
Rank ID Comments Likes/Claps Author Affiliation Venue
1 G12 2 348 Kathan Shah & Subhra Mondal Lyft Medium
2 G11 - 156 Prateek Sachan National Australia Bank (NAB) Medium
3 G15 - 128 Nikos Katirzis Expedia Group Medium
4 G25 - 52 Russ Miles Chaos Toolkit Medium
5 G14 1 24 Mathew Moon Just Eat Takeaway.com Medium
Figure 11. Research Type of Academic (AL) and Grey (GL) Literature Studies.
Figure 12. Venue Ranking of Academic Literature Studies.
7. Discussion
In this section, we summarize the key findings of our review and discuss their implications. We also outline open issues in chaos engineering research and discuss potential threats to our study’s validity.
7.1. Implications of Findings
The analysis highlights five key implications for researchers and practitioners in chaos engineering:
  • Need for a Unified Definition: The literature reveals inconsistencies in defining chaos engineering. Although tech giants like Google, Amazon, and Netflix provide definitions that emphasize resilience testing through controlled failure injection, there is variation in focus across other companies, with some prioritizing production-level experimentation, self-healing capabilities, or anti-fragility. A unified definition can align research efforts and support the systematic adoption of chaos engineering practices, strengthening reliability across complex Information Technology (IT) environments (Siswanto et al., 2024; Jayakody and Wijayanayake, 2024; Jones, 2019).
  • Complementary Role to Traditional Testing: Chaos engineering should be viewed as a complement to traditional testing rather than a replacement (Bocetta, 2019; Palani and Gupta, 2023). While conventional testing may miss subtle vulnerabilities within complex systems, chaos engineering exposes these issues through proactive experimentation. Recognizing chaos engineering as an additional layer to traditional methods emphasizes its role in enhancing system resilience, especially in high-stakes, mission-critical environments. Integrating chaos engineering into testing workflows is encouraged for those aiming to build more robust and reliable systems.
  • Growing Alignment of Research Theory and Practice: The prevalence of solution-oriented papers in academic literature alongside experience reports in grey literature reflects an increasing alignment between research and practice in chaos engineering. This synergy allows researchers to refine models based on real-world applications while practitioners benefit from research-backed frameworks, thus bridging the gap between theory and implementation. As suggested in previous studies (Jernberg et al., 2020; Rosenthal and Jones, 2020), fostering industry-academia collaborations could further integrate academic rigor with practical insights, advancing chaos engineering as a more resilient and impactful field.
  • Targeted Tool Selection: The study suggests that organizing chaos engineering tools by factors like execution environment and automation level might help practitioners make informed decisions when selecting tools. This taxonomy enables organizations to choose tools that match their specific infrastructure, operational needs, and risk tolerance, minimizing disruptions, managing costs, and enhancing resilience testing across development stages. This targeted approach allows tools to align with specific testing phases, from initial experimentation to production-level assessments, ensuring that chaos engineering efforts are both cost-efficient and impactful, ultimately supporting system reliability.
  • Standardized Practices: The findings highlight the importance of a standardized approach in chaos engineering to ensure consistent and practical application across diverse environments. Following best practices, which contribute to reliable and predictable outcomes (Kumara et al., 2021), enables organizations to build resilience and systematically manage risks. These practices, such as starting with small, controlled experiments, setting clear objectives, and automating processes where possible, support organizations in achieving reliable outcomes crucial for managing complex systems. Conversely, neglecting these practices, such as misaligned experiment design or insufficient safety mechanisms, can lead to inconsistent testing and increased system vulnerability. These findings underscore the need to balance adherence to best practices with avoiding common pitfalls, ensuring that chaos engineering aligns with broader operational goals to foster stability and reliability.
  • Insights through Quantitative and Qualitative Fusion: An essential implication of the findings is that combining quantitative and qualitative evaluation methods provides a comprehensive framework for assessing and improving system stability in chaos engineering. Quantitative metrics offer objective data on performance under stress, such as error rates, MTTR, and availability. At the same time, qualitative insights, including user experience and team feedback, reveal the broader impacts of system failures. This combined approach helps organizations identify technical vulnerabilities and understand their user and business impacts, enabling targeted investments in both stability and operational efficiency to align improvements with business goals and user expectations.
7.2. Open Research Issues and Practical Challenges in Chaos Engineering
Our review identifies several unresolved issues in chaos engineering, including research gaps and organizational challenges. We examine these issues from three perspectives: people, process, and technology.
7.2.1. People
In chaos engineering, people include engineers, managers, and other team members who collaborate to design and run experiments, ensuring that the systems are thoroughly tested (Moon, 2022; Nikolaidis et al., 2021; Graglia, 2024). First, some organizations struggle with poor communication and collaboration between teams, leading to incomplete or ineffective experiments  (Zhang et al., 2019). Encouraging more vital teamwork and shared ownership could improve results by fostering better department communication (Gill, 2021; Katirtzis, 2022; Kostic, 2024). Another challenge is the shortage of specialists, making it difficult for organizations to carry out chaos experiments. This gap could be addressed by creating comprehensive training programs and certifications that allow staff to develop the necessary expertise (Dedousis et al., 2023; Nikolaidis et al., 2021; Yu et al., 2021). Finally, cultural resistance is expected, as fear of failure and blame often hinder wider adoption  (Simonsson et al., 2021; Rosenthal and Jones, 2020). Implementing cultural change initiatives that promote learning from controlled failures and reduce fear of backlash can create a more open environment for experimentation (Green, 2023).
7.2.2. Process
The process aspect refers to the specific steps and methodologies, such as planning, testing, and reporting, that guide the implementation of chaos experiments (Bairyev, 2023; Kostic, 2024; Olmstead, 2023). First, the lack of standardized guidelines across industries makes consistent application difficult (Nikolaidis et al., 2021; Cotroneo et al., 2022; Kostic, 2024). Developing a set of best practices and frameworks would help organizations measure progress and improve the effectiveness of these practices (Shah, 2021; Bremmers, 2021). The research shows that many organizations struggle to prioritize chaos experiments alongside daily operations, often finding it difficult to allocate time and resources while maintaining operational stability. Establishing transparent processes for scheduling and running these experiments during non-critical periods or as part of routine maintenance can help mitigate the issue (Kesim et al., 2020).
7.2.3. Technology
Technology refers to the systems and infrastructure used to design, execute, and monitor failure scenarios, supporting and amplifying human efforts and processes (Graglia, 2024; Olmstead, 2023). First, chaos experiments are resource-intensive and require significant computational power and storage. This makes it difficult for organizations to allocate infrastructure without affecting critical operations  (Dedousis et al., 2023; Nikolaidis et al., 2021). To manage this, companies can schedule experiments during low-traffic periods or use scalable cloud resources to reduce strain  (Kostic, 2024; Gill, 2021). Second, deploying chaos tools requires specialized knowledge, making it challenging to scale experiments consistently (Zhang et al., 2019; Yu et al., 2021). This challenge can be alleviated with better documentation, training, and platforms that simplify deployment for teams with less expertise (Bairyev, 2023). Third, many chaos engineering tools lack automation, making setup and execution time-consuming and prone to errors. Automating processes like fault injection and monitoring can reduce the burden on engineers, minimize mistakes, and improve scalability. Organizations can address this through commercial tools or internal solutions (Bairyev, 2023; Cotroneo et al., 2022; Shah, 2021). Finally, while Artificial Intelligence (AI) holds promise for further streamlining experiment design and execution, research in this area is still limited. Using AI could reduce errors, improve efficiency, and help organizations scale their resilience efforts more effectively (Zhang et al., 2023).
7.3. Threats to Validity
We considered several potential threats to the external, construct, internal, and conclusion validity (Wohlin et al., 2012) that may impact our study.
7.3.1. Threats to External Validity
External validity refers to the extent to which the study results are applicable in a broader context (Wohlin et al., 2012). Since our multivocal review incorporated studies from various online sources, our findings may have limited applicability across the wider landscape of practices and general disciplines in chaos engineering, thus posing a threat to external validity. To enhance the external validity of our findings, we conducted multiple feedback sessions throughout the analysis of the selected studies. The discussions of each session were analyzed, providing qualitative insights that allowed us to refine our research methods and the applicability of our results. Furthermore, we compiled an online appendix with all artifacts generated during our analysis, including a complete list of sources, codes, and groups (see Section  2), making them publicly accessible for further examination. Another potential threat is the risk of missing relevant literature (grey and scientific) due to alternative terminology for related concepts that may not match our search terms. We explicitly included relevant associated synonyms and terms in our search strings to address this threat. We also leveraged search engine features that inherently consider variations of terms in a search string. The items identified were rigorously assessed based on multiple quality dimensions (Wohlin, 2014). Finally, while there are numerous chaos engineering tools, we limited our analysis to ten due to practicality. To mitigate this limitation, we selected the ten most relevant tools currently used by practitioners. A detailed comparison of these chosen tools may provide additional insight, which we propose for future research.
7.3.2. Threats to Construct and Internal Validity
Construct validity pertains to the generalizability of the study’s constructs, while internal validity relates to the credibility of the methods used for data analysis, including possible biases (Wohlin et al., 2012). To reduce potential threats to construct and internal validity, we implemented multiple triangulation rounds, inter-rater reliability assessments, and quality control measures (see Section 2). These methods helped minimize potential biases, such as observer and interpretive biases, thereby enhancing the validity of our data analysis. Additional studies and codes were incorporated, although no further inter-rater evaluation was conducted. In particular, some risk of observer bias remains inherent in this methodology (Kumara et al., 2021).
7.3.3. Threats to Conclusion Validity
Conclusion validity refers to the extent to which the conclusions of the study are credibly supported by the available data  (Wohlin et al., 2012). To address this threat, we used thematic coding and inter-rater reliability assessments to reduce observer and interpretive biases, aiming to analyze the data thoroughly. Furthermore, the study conclusions were independently derived by different authors and subsequently reviewed and validated against the primary studies selected during a feedback session.
8. Related Surveys
In this section, we consider key surveys on resilience, robustness, and fault injection techniques, exploring their relationship to the proactive focus of chaos engineering. For example, Natella et al. conducted an extensive survey on Software Fault Injection (SFI) techniques, categorizing methods by fault types, injection approaches, and application domains. These techniques are valuable for assessing software reliability in controlled settings. However, SFI techniques are generally applied in pre-production or isolated environments rather than in the live, iterative fault-testing scenarios central to chaos engineering (Natella et al., 2016). Another study, conducted by Colman-Meixner et al. (Colman-Meixner et al., 2016), explores resilience strategies across cloud infrastructure layers, such as physical servers, virtualization, and application components. Their work proposes a multi-layered framework with redundancy and failover mechanisms to ensure continuity during failures. While this approach aligns with the resilience goals of chaos engineering, the focus remains on structured recovery strategies within controlled environments rather than proactive fault injection in live settings. Laranjeiro et al. (Laranjeiro et al., 2021) reviewed robustness assessment techniques for systems such as operating systems and middleware and discussed methods such as fault injection and API testing to evaluate dependability under stress. These approaches aim to identify weaknesses but emphasize static testing, which does not include the proactive, live-system experiments that chaos engineering relies on. More recently, Rouholamini et al. (Rouholamini et al., 2020) presented a study on proactive self-healing techniques in cloud computing, categorizing strategies such as automated fault detection, anomaly recognition, and self-correction mechanisms. These techniques improve system resilience by addressing faults as they arise, but differ from the chaos engineering approach of using intentional fault injection to reveal latent vulnerabilities preemptively.
In general, there is a significant knowledge gap in the area of proactive and iterative resilience testing directly within live production environments. Moreover, there are no comprehensive reviews of the literature on chaos engineering. Therefore, this review paper systematically analyzes, assesses, and summarizes the existing literature to clarify the unique contributions of chaos engineering to the domain of system resilience.
9. Conclusion and Future Work
As modern distributed systems become increasingly complex, chaos engineering offers a proactive approach to ensure stability by introducing controlled failures that reveal system weaknesses, prevent outages, and support high performance under stress. This study bridges the gap between research insight and industry practices through a multivocal literature review (MLR) of 88 sources from academic and grey literature. To promote clarity across academia and industry, we proposed a unified definition of chaos engineering, reinforcing its role as a vital complement to traditional testing. We identified the main functionalities of a chaos engineering platform and compiled a set of principles that can guide systematically planning and conducting chaos experiments. In addition, we presented the key components of a chaos engineering platform and identified 12 quality requirements to guide the selection and evaluation of the platform. We also identified eight qualitative and four quantitative metrics essential for evaluating the impact of chaos experiments, ensuring that chaos engineering meaningfully enhances system robustness. Our review highlights four technical and three socio-technical challenges that chaos engineering seeks to resolve. To support organizations in structuring chaos engineering practices, we developed a taxonomy that organizes tools and techniques by environment, automation mode, automation strategy, and deployment stage while identifying the ten most commonly used tools within this framework. This categorization is valuable to practitioners and researchers when selecting tools that align with specific infrastructure needs and risk profiles. In addition, we identified six best practices and five common pitfalls in the implementation of chaos engineering. These practices and potential pitfalls guide organizations in optimizing their chaos engineering initiatives to foster safer and more robust systems. Our findings also reveal that solution-oriented research in academia and practical experience in industry converge, underscoring chaos engineering’s essential role in strengthening modern distributed systems. Finally, we outlined several research and practical challenges that must be addressed to realize chaos engineering’s vision entirely. We believe that these identified open issues can guide future research in the domain of chaos engineering.
In future work, we plan to use GitHub repository mining to identify various open-source projects that implement chaos engineering practices and utilize associated tools. This exploration will complement the ten tools analyzed in this literature review, offering broader insights into the prevalence and practical application of chaos engineering in diverse projects. Additionally, we aim to apply chaos engineering to improve the robustness of AI-enabled systems by evaluating how such systems handle disruptions in accuracy, adaptability, and consistency, ensuring that they continue to deliver reliable performance even under challenging conditions.